### v1

# On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation

Abstract
This paper proposes a maintenance strategy for a satellite constellation that utilizes on-orbit servicing (OOS). Under this strategy, the constellation operator addresses satellite failures in two ways: by deploying new satellites and by recovering failed satellites through OOS. We develop an inventory management model with a parametric replenishment policy for the maintenance process, which can evaluate the performance of the satellite constellation system. Based on this model, we formulate two single-objective optimization problems representing the decision-making contexts of two main stakeholders—the constellation operator and the OOS provider—and a bi-objective optimization problem that can reflect the tension between the two. A case study of the OOS-supported maintenance for a real-world-scale constellation provides valuable insights that help explain the behaviors of the stakeholders.
Nomenclature
$\mathrm{AMC}$  = constellation operator’s annual maintenance cost, \$M/year
$\mathrm{AP}$  = OOS provider’s annual profits, \$M/year
$C_{\mathrm{hold}}$  = annual holding cost, \$M/year
$C_{\mathrm{lau}}$  = annual launch cost, \$M/year
$C_{\mathrm{maneuv}}$  = annual maneuvering cost, \$M/year
$C_{\mathrm{manufac}}$  = annual manufacturing cost, \$M/year
$C_{\mathrm{oos}}$  = annual OOS cost, \$M/year
$c_{\mathrm{hold}}$  = annual holding cost of a satellite, \$M/satellite/year
$c_{\mathrm{lau}}$  = launch costs, \$M
$c_{\mathrm{manufac}}$  = manufacturing cost of a satellite, \$M/satellite
$c_{\mathrm{oos}}$  = cost for the OOS provider to recover a satellite, \$M
$\mathrm{ES}_{\mathrm{parking}}$  = expected shortage of parking spares, -
$\mathrm{ES}_{\mathrm{plane}}$  = expected shortage of in-plane spares, in units of satellites
$h_{\mathrm{parking}}$  = altitude of parking orbits, km
$h_{\mathrm{plane}}$  = altitude of orbital planes in constellation, km
$i$  = inclination angle of an orbit, deg
$k_{s}$  = reorder point of parking spares, -
$k_{Q}$  = order quantity of parking spares, -
$M_{\mathrm{dry}}$  = dry mass of satellite, kg
$M_{\mathrm{fuel}}$  = fuel mass required for orbital transfer, kg
$\dot{M}_{\mathrm{prop}}$  = mass flow rate of satellite propulsion system, kg/s
$N_{\mathrm{oos}}$  = the maximum number of service for a satellite, -
$N_{\mathrm{parking}}$  = number of parking orbits, in units of planes
$N_{\mathrm{plane}}$  = number of orbital planes in constellation, in units of planes
$N_{\mathrm{sat}}$  = number of satellites per orbital plane in constellation, in units of satellites
$N_{t}$  = number of time units per year, in units of time per year
$n_{\mathrm{oos}}$  = annual service frequency, per year
$n_{\mathrm{parking}}$  = annual order frequency of parking spares, per year
$n_{\mathrm{plane}}$  = annual order frequency of in-plane spares, per year
$p_{\mathrm{oos}}$  = OOS price, \$M
$Q$  = order quantity of in-plane spares, in units of satellites
$Q_{\max}$  = launch capacity, in units of satellites
$\overline{\mathrm{SL}}_{\mathrm{parking}}$  = mean stock level of parking spares, -
$\overline{\mathrm{SL}}_{\mathrm{plane}}$  = mean stock level of in-plane spares, in units of satellites
$\overline{\mathrm{SL}}_{\mathrm{wait}}$  = mean stock level of failed satellites waiting for service, in units of satellites
$r_{\mathrm{oos}}$  = fraction of serviceable in-orbit failures, - (or %)
$s$  = reorder point of in-plane spares, in units of satellites
$T_{d}$  = mean time to disposal of satellites serviced up to the maximum, in time units
$T_{\mathrm{life}}$  = maximum lifespan of satellites, years
$T_{\mathrm{oos}}$  = mean time spent during awaiting OOS until disposal, in time units
$T_{\mathrm{opr}}$  = mean time spent during operations until disposal, in time units
$T_{\mathrm{parking}}$  = mean time spent as parking spares until disposal, in time units
$T_{\mathrm{plane}}$  = mean time spent as in-plane spares until disposal, in time units
$T_{\mathrm{trans}}$  = mean time spent during orbital transfer from parking to target orbit, in time units
$\mathrm{TOF}$  = time of flight, in time units
$t_{\mathrm{trans}}$  = in-plane transfer maneuver time, in units of time
$\mathcal{T}_{j}$  = admissible time interval for supplementation from the jth closest parking orbit
$V_{\mathrm{e}}$  = exhaust velocity of satellite propulsion system, km/s
$\mathbf{x}_{\mathrm{con}}$  = decision variables vector of constellation operator
$\mathbf{x}_{\mathrm{oos}}$  = decision variables vector of OOS
$\alpha_{1},\alpha_{2}$  = shape parameters of cost-responsiveness function, -
$\gamma_{m}$  = fraction of in-plane spares serviced $m$ times, - (or %)
$\delta$  = random variable for stochastic demand
$\epsilon_{\mathrm{fuel}}$  = fuel mass conversion coefficient, \$M/kg
$\lambda_{\mathrm{parking}}$  = demand rate at parking spares, per unit time
$\lambda_{\mathrm{parking}}$  = demand rate at in-plane spares, in units of satellites per unit time
$\lambda_{\mathrm{sat}}$  = failure rate of a satellite, in units of satellites per year
$\mu_{\mathrm{lau}}$  = parameter of stochastic launch delay, in units of inverse time
$\mu_{\mathrm{oos}}$  = responsiveness of service, in units of inverse time
$\rho_{\mathrm{parking}}$  = order fill rate for parking spares, - (or %)
$\rho_{\mathrm{plane}}$  = order fill rate for in-plane spares, - (or %)
$\tau$  = random variable for stochastic lead time
$\mathbb{R}_{++}$  = positive real number set
$\mathbb{Z}_{+}$, $\mathbb{Z}_{++}$  = non-negative, and positive integer sets
1 Introduction
Satellite mega-constellations have played a pivotal role in shaping the modern space industry, driven by the demand for global coverage services and advances in space technologies. These driving forces have encouraged active participation from major players, including SpaceX’s Starlink, Amazon’s Project Kuiper, and Eutelsat’s OneWeb [1, 2]. This participation has expanded the market and established mega-constellations as a critical component in the ecosystem of the space industry.
From a system maintenance perspective, satellite mega-constellations expose operators to more frequent component failures (i.e., satellite failures), primarily due to the reduced reliability of individual satellites and the vastly increased number of units in operation. Addressing these challenges necessitates a systematic maintenance strategy, specifically, a well-structured logistics framework for effective failure response [3]. Recent studies, including those by Jakob et al. [4] and Kim et al. [5], have proposed strategies and corresponding assessment tools to tackle this problem.
Specifically, Jakob et al. [4] proposed a multi-echelon supply chain for spare satellites, along with an accompanying inventory management policy and analytical forms of performance metrics. The supply chain comprises three echelons—ground spares, parking spares, and in-orbit spares—which are analogous to suppliers, warehouses, and retailers, respectively, in a terrestrial supply chain. Within this framework, parking spares and in-plane spares are replenished using distinct $\left(s,Q\right)$ policies [6], where a batch of $Q$ units is ordered once the stock level drops to $s$. Under this structure, the authors assumed a Poisson distribution for satellite failures and an exponential distribution for launch lead times. They derived analytical expressions for inventory-related parameters, maintenance costs, and service levels. The cost model encompasses several elements, including satellite manufacturing, launch, holding, and maneuvering. Finally, they formulated an optimization problem from the constellation operator’s perspective to determine the policy parameters and parking orbit geometries that minimize maintenance costs while satisfying the required service level of the maintenance system.
Building on the work of Jakob et al., Kim et al. [5] established a dual supply strategy by introducing an auxiliary (direct) channel that directly injects spare satellites into the target orbits [7]. They integrated an inventory policy denoted by $\left(R_{1},R_{2},Q_{1},Q_{2}\right)$ with a time window for in-plane spares. Under this policy, when the stock level falls to $R_{1}$, $Q_{1}$ units are ordered from the parking spares via the primary (indirect) supply channel. If the stock level subsequently drops to $R_{2}$ within the predefined time window, the auxiliary (direct) supply channel is activated, and $Q_{2}$ units are ordered. They derived analytical expressions for inventory-related parameters and formulated two optimization problems from different perspectives: the satellite constellation operator and the auxiliary launch service provider (i.e., the transportation service provider for the direct channel). For a comprehensive review of maintenance for small satellite constellations, readers are referred to Jakob et al. [4] and Kim et al. [5].
Beyond deploying spare satellites, on-orbit servicing (OOS) can recover a satellite’s operational capability and is beginning to be commercialized [8, 9, 10, 11]. Historically, OOS has targeted heavy space systems, such as the International Space Station, the Hubble Space Telescope, and communication satellites at geosynchronous equatorial orbit, whose life extension offers significant monetary value. Numerous studies have quantified OOS’s benefits for single large systems [12, 13, 14, 15, 16, 17, 18], and operational successes, such as the MEV-1 mission [19], underscore its practical value.
More recently, researchers have begun to consider OOS for low Earth orbit (LEO) constellations, where the rapid growth of space activities presents new opportunities. Although each individual satellite is relatively inexpensive and the added value of servicing a single unit is smaller compared to larger space systems, the overall scale of constellations continues to attract players in the space industry. Luu and Hastings [20] compared two strategies—deploying spare satellites and using OOS—for maintaining satellite mega-constellations. However, existing research has not yet examined spare satellites and OOS simultaneously as part of a unified maintenance strategy for satellite mega-constellations.
To address this gap, we propose a novel maintenance strategy for satellite constellations that integrates OOS with the deployment of spare satellites. Within this strategy, failed satellites are not only replaced by spare satellites but can also be repaired through OOS if the failure is recoverable. To evaluate the performance of the proposed strategy at a strategic level, we construct the corresponding supply chain and develop an assessment tool based on an inventory management model. Building on this foundation, we introduce relevant decision-making scenarios, formulate them as optimization problems, and conduct a case study to demonstrate the applicability of the proposed approach and provide actionable insights.
The remainder of this paper is organized into five sections. Section 2 introduces the modeling assumptions and constructs the supply chain and inventory management framework for evaluating the OOS-integrated maintenance strategy. Section 3 describes the decision-making scenarios, while Section 4 presents a case study that illustrates the application of these scenarios. Finally, Section 5 concludes the paper.
2 Model Construction
This section outlines the proposed maintenance strategy and the evaluation framework. We begin by introducing the supply chain and the underlying assumptions. Next, we present the parametric models that describe the mean behavior of the inventory at each echelon of the supply chain and construct the maintenance cost and system performance parameters. Finally, we assess the model’s accuracy by comparing its results with those obtained from discrete-time simulations.
2.1 Supply Chain Overview and Assumptions
We assume that the target system for maintenance has a Walker Delta pattern [21], with inclination $i$ and altitude $h_{\mathrm{plane}}$. It consists of $N_{\mathrm{plane}}$ orbital planes whose right ascensions of ascending nodes (RAANs) are equally spaced by $2\pi/N_{\mathrm{plane}}$. Each orbital plane contains $N_{\mathrm{sat}}$ operational satellites.
The operator sources spare satellites in two ways: 1) recovering the operational life of failed satellites via OOS, and 2) launching newly manufactured satellites from the ground. Figure 1 presents an overview of the proposed supply chain.
Figure 1: Overview of the supply chain for the OOS-integrated maintenance strategy.
2.1.1 Sourcing Method 1: Life Extension by OOS
OOS addresses serviceable failures. The specific failure modes that are considered serviceable depend on the spacecraft design (e.g., service-friendly interfaces, degree of modularity), the maturity of OOS technologies, and the availability of supporting infrastructure. At the strategic level, we assume a fixed value for the fraction of serviceable in-orbit failures, $r_{\mathrm{oos}}$. In other words, if 100 failures occur, then on average $100r_{\mathrm{oos}}$ are serviceable and $100\left(1-r_{\mathrm{oos}}\right)$ are not. When a serviceable failure occurs, the failed satellite waits for servicing. After servicing, it is assumed to fully regain its operational capability and remain on standby as an in-plane spare.
In addition, even if a serviceable failure occurs, the operator may choose not to recover the satellite if other components, those that OOS cannot renew, are obsolete and require replacement. Therefore, we further assume that each satellite has an operational lifespan, denoted by $T_{\mathrm{life}}$. For example, if OOS can only refuel a depleted satellite, other components (e.g., electronics, structure) remain unchanged. In such cases, repeatedly refueling the satellite beyond its structural or technological lifespan may be inadequate. Consequently, during the strategic phase, the operator may define $T_{\mathrm{life}}$ based on the OOS capabilities and spacecraft design. As a proxy for $T_{\mathrm{life}}$, we introduce the maximum number of OOS per satellite, $N_{\mathrm{oos}}$, as a tracking parameter.
The service lead time, defined as the duration between service request reception and service completion, is primarily composed of a series of maneuvers performed by the servicing spacecraft to approach the target satellite (rendezvous, proximity operations, docking), along with the time required to execute the servicing activity itself. The lead time depends on the current state of the OOS provision system. For example, if an available servicing spacecraft is spatially distant from the target satellite at the time of the service request, the lead time will be long. Since the time required for the available servicing spacecraft to maneuver to the target is inherently random, the lead time is modeled as a probabilistic variable.
2.1.2 Sourcing Method 2: Deploying New Satellites
If a failure is non-serviceable or the failed satellite has exceeded its lifespan, a spare satellite replaces the failed one. They are managed by the strategy proposed by Jakob et al. [4], which is the indirect channel in [5]. Spare satellites are distributed across three locations—ground facilities (ground spares), parking orbits (parking spares), and target orbits (in-plane spares). Here, the parking orbits share the same inclination as the target orbits, $i$, but have a lower altitude, $h_{\mathrm{parking}}<h_{\mathrm{plane}}$. There are $N_{\mathrm{parking}}$ parking planes, with RAANs equally spaced by $2\pi/N_{\mathrm{parking}}$.
In-plane spares directly replace failed satellites and are distributed across the (near-)target orbits (i.e., the operational planes of the constellation). When the inventory of in-plane spares reaches a reorder point $s$, a batch of $Q$ spares is sourced from the parking spares, the next echelon. Here, the spare satellites are transported by leveraging $J_{2}$ perturbation to align RAANs of the parking and target orbits, and the satellite’s propulsion system to raise the altitude. Here, we have the RAAN drift rate relation [22]:
$\dot{\Omega}\left(a,e,i\right)=-\frac{3}{2}\sqrt{\frac{\mu_{E}}{a^{3}}}\frac{R_{E}^{2}}{a^{2}(1-e^{2})^{2}}J_{2}\cos{i},$ (1)
where $a$, $i$, and $e$ are the semi-major axis, the inclination, and the eccentricity of the orbit, respectively, $J_{2}$ is the zonal harmonic coefficient, $R_{E}$ is the Earth’s equatorial radius, and $\mu_{E}$ represents the Earth’s gravitational constant.
Parking spares perform orbit-raising maneuvers to replenish the in-plane spares once the RAANs of the corresponding parking and target orbits are aligned. The motion is modeled using Keplerian dynamics under continuous thrust, reflecting the increasing adoption of low-thrust propulsion systems—such as electric thrusters—by small satellites due to their high efficiency. The required velocity increment ($\Delta V$) and time of flight ($\mathrm{TOF}$) are as follows [23, 24]:
$\Delta V=\sqrt{\frac{\mu_{E}}{r_{\mathrm{initial}}}}-\sqrt{\frac{\mu_{E}}{r_{\mathrm{final}}}},$ (2)
$\mathrm{TOF}=\frac{M_{\mathrm{fuel}}}{\dot{M}_{\mathrm{prop}}},$ (3)
where $r_{\mathrm{initial}}$ and $r_{\mathrm{final}}$ are the initial and final orbit radii, respectively ($r_{\mathrm{initial}}<r_{\mathrm{final}}$), $\mu_{E}$ is Earth’s gravitational constant, $M_{\mathrm{fuel}}$ is the required fuel mass, and $\dot{M}_{\mathrm{prop}}$ is the mass flow rate of the propulsion system. Here, the fuel mass follows the rocket equation:
$M_{\mathrm{fuel}}=M_{\mathrm{dry}}(e^{\Delta V/V_{e}}-1),$ (4)
where $M_{\mathrm{dry}}$ is the satellite’s dry mass and $V_{e}$ is the exhaust velocity.
Parking spares are managed in batches of size $Q$ to match the in-plane replenishment batch size. When the inventory of parking spares falls to $k_{s}$ batches, then $k_{Q}$ batches are newly launched from the ground to the parking orbit. Ground spares are assumed to be always available.
We construct an inventory management model based on the assumptions listed below:
  1. 1.
    In-orbit failures of operational satellites follow a Poisson process with constant rate $\lambda_{\mathrm{sat}}$.
  2. 2.
    Spare satellites’ failures are neglected (i.e., ground, parking, and in‐plane spares are assumed failure-free).
  3. 3.
    In-plane spares are available instantaneously to replace a failed satellite [4].
  4. 4.
    Each target orbit is replenished from the closest available parking orbit (i.e., a nonempty parking orbit with the earliest RAAN alignment). When an in-plane stock level hits its reorder point ($s$), a batch of size $Q$ is drawn from that parking orbit [4].
  5. 5.
    Ground-to-parking delivery can supply only a single parking orbit [4].
  6. 6.
    A new in-plane order is placed only after the previous one has completed transit to facilitate the tracking of the orders [4].
  7. 7.
    Stock-outs are rare; therefore, the probability of all parking orbits being out of stock is negligible [4].
  8. 8.
    Both the reorder point and order quantity at parking orbits are constrained to integer multiples of the batch size ($Q$) [4].
  9. 9.
    An OOS lead time follows an exponential distribution, which is a stochastic distribution characterized by a single parameter, with mean time to recovery (or equivalently, mean time to repair, MTTR), $\mu_{\mathrm{oos}}^{-1}$.
2.2 In-Plane Spares
In-plane spares directly respond to failures of operational satellites, which follow a Poisson process with rate parameter $\lambda_{\mathrm{sat}}$. The failure rate within a constellation’s operational plane is defined as
$\lambda_{\mathrm{plane}}=\frac{\lambda_{\mathrm{sat}}N_{\mathrm{sat}}}{N_{t}},$ (5)
where $N_{\mathrm{sat}}$ is the number of satellites per plane, and $N_{t}$ is the number of discretized time units per year [4].
In-plane spares are replenished by parking spares and OOS. Figure 2 provides a closer view of the closed-loop system shown in Fig. 1, illustrating the flows of spare and failed satellites for the case of $N_{\mathrm{oos}}=2$. Since our model represents the mean behavior over an infinite time horizon, we define $\gamma_{0}$ as the fraction of newly arrived in-plane spares sourced from parking orbits, and $\gamma_{m}$ ($m=1,2,\dots,N_{\mathrm{oos}}$) as the fraction of in-plane spares that have been serviced $m$ times, also in the infinite time horizon.
Figure 2: Example spare/failed satellites flows within a target orbit, $N_{\mathrm{oos}}=2$.
Assuming that the service lead time follows an exponential distribution with the mean value of $\mu_{\mathrm{oos}}^{-1}$ implies that the OOS process recovers a failed satellite with probability $\mu_{\mathrm{oos}}$ per unit time [25]. Based on this assumption, the flow balance equations at the steady-state for failed satellites awaiting OOS and in-plane spares over an infinite horizon are as follows:
$\lambda_{\mathrm{plane}}\gamma_{m-1}r_{\mathrm{oos}}=\overline{\mathrm{SL}}_{\mathrm{wait},m}\mu_{\mathrm{oos}},\text{ for }m=1,2,\dots,N_{\mathrm{oos}},$ (6)
$\overline{\mathrm{SL}}_{\mathrm{wait},m}\mu_{\mathrm{oos}}=\lambda_{\mathrm{plane}}\gamma_{m},\text{ for }m=1,2,\dots,N_{\mathrm{oos}},$ (7)
where $\overline{\mathrm{SL}}_{\mathrm{wait},m}$ is the mean stock level of failed satellites that have been serviced $m$ times and are awaiting the next service. Also, the fractions $\gamma_{m}$ satisfy the normalization condition:
$\sum_{m=0}^{N_{\mathrm{oos}}}\gamma_{m}=1.$ (8)
Solving these balance equations determine the fractions $\gamma_{m}$ ($m=0,1,\dots,N_{\mathrm{oos}}$) as follows:
$\gamma_{0}=\frac{1-r_{\mathrm{oos}}}{1-\left(r_{\mathrm{oos}}\right)^{N_{\mathrm{oos}}+1}},$ (9)
$\gamma_{m}=r_{\mathrm{oos}}\gamma_{m-1},\text{ for }m=1,2,\dots,N_{\mathrm{oos}}.$ (10)
In addition, we define $\overline{\mathrm{SL}}_{\mathrm{wait}}$ as follows:
$\overline{\mathrm{SL}}_{\mathrm{wait}}=\sum_{m=1}^{N_{\mathrm{oos}}}\overline{\mathrm{SL}}_{\mathrm{wait},m}.$ (11)
The demand and recovery processes are modeled as Poisson processes. Since the stock drop corresponds to demand minus recovery, the resulting stock drop process follows a Skellam distribution [26]. Accordingly, the probability mass function (PMF) of the stock drop over a time interval $\tau$ is given by:
$P_{\mathrm{plane}}\left(\delta;\tau\right)=e^{-\lambda_{\mathrm{plane}}\tau\left(2-\gamma_{0}\right)}\frac{1}{\left(1-\gamma_{0}\right)^{\delta/2}}I_{\delta}\left(2\lambda_{\mathrm{plane}}\tau\sqrt{1-\gamma_{0}}\right),\quad\delta\in\mathbb{Z}_{+},$ (12)
where $I_{\delta}$ denotes the modified Bessel function of the first kind [27].
To respond to this stock level drop, we adopt an $\left(s,Q\right)$ replenishment policy: whenever the in-plane spare inventory level drops to the reorder point $s$, an order of size $Q$ is placed. The lead time for such an order depends on the availability of parking spares, and the time for RAAN alignment between the target and parking orbits and for orbit raising [4, 5]. Denoting the lead time by $\tau$, its probability density function (PDF) is given as:
$f_{\mathrm{plane}}\left(\tau\right)=P^{\prime}_{\mathrm{av}}\left(j\right)\frac{N_{\mathrm{parking}}\left|\dot{\Omega}_{\mathrm{rel}}\right|}{2\pi},\text{ for }\tau\in\mathcal{T}_{j},\ j=1,2,\dots,N_{\mathrm{parking}},$ (13)
where $P_{\mathrm{av}}^{\prime}\left(j\right)$ is the normalized probability that the $j$th closest parking orbit is the closest available parking one, $\dot{\Omega}_{\mathrm{rel}}$ is the relative RAAN drift rate between target and parking orbits, and $\mathcal{T}_{j}$ is the admissible range of alignment time between a target orbit and its $j$th closest parking orbit, defined as:
$P_{\mathrm{av}}^{\prime}\left(j\right)=\frac{P_{\mathrm{av}}\left(j\right)}{\sum_{j^{\prime}=1}^{N_{\mathrm{parking}}}P_{\mathrm{av}}\left(j^{\prime}\right)},\quad P_{\mathrm{av}}\left(j\right)=\left(1-\rho_{\mathrm{parking}}\right)^{j-1}\rho_{\mathrm{parking}},\text{ for }j=1,2,\dots,N_{\mathrm{parking}},$ (14)
$\dot{\Omega}_{\mathrm{rel}}=\dot{\Omega}\left(h_{\mathrm{parking}},0,i\right)-\dot{\Omega}\left(h_{\mathrm{plane}},0,i\right),$ (15)
$\mathcal{T}_{j}=\left[\left(j-1\right)\frac{2\pi}{h_{\mathrm{parking}}}\frac{1}{\left|\dot{\Omega}_{\mathrm{rel}}\right|}+t_{\mathrm{trans}},j\frac{2\pi}{N_{\mathrm{parking}}}\frac{1}{\left|\dot{\Omega}_{\mathrm{rel}}\right|}+t_{\mathrm{trans}}\right),\text{ for }j=1,2,\dots,N_{\mathrm{parking}}.$ (16)
Here, $t_{\mathrm{trans}}$ is the time of flight from a parking orbit to a target orbit calculated by Eq. (3).
Following these demand and supply models, we define several inventory parameters. The mean stock level ($\overline{\mathrm{SL}}_{\mathrm{plane}}$) represents the average inventory held. The order frequency ($n_{\mathrm{plane}}$) denotes how often orders are placed. The expected shortage ($\mathrm{ES}_{\mathrm{plane}}$) is the average shortage within a reorder cycle, and the order fill rate ($\rho_{\mathrm{plane}}$) is the complement of the ratio between expected shortage and order quantity, representing the service level of the maintenance system. The service frequency ($n_{\mathrm{oos}}$) indicates the frequency of service calls. These are expressed as follows [4, 5, 6]:
$\overline{\mathrm{SL}}_{\mathrm{plane}}=\int_{\tau\geq t_{\mathrm{trans}}}\left(s-\lambda_{\mathrm{plane}}\gamma_{0}\tau+\frac{Q}{2}+\frac{1}{2}\right)f_{\mathrm{plane}}\left(\tau\right)\,d\tau,$ (17)
$n_{\mathrm{plane}}=\frac{\lambda_{\mathrm{plane}}\gamma_{0}}{Q}N_{t},$ (18)
$\mathrm{ES}_{\mathrm{plane}}=\int_{\tau\geq t_{\mathrm{trans}}}\sum_{\delta\geq s}\left(\delta-s\right)P_{\mathrm{plane}}\left(\delta;\tau\right)f_{\mathrm{plane}}\left(\tau\right)\,d\tau,$ (19)
$\rho_{\mathrm{plane}}=1-\frac{\mathrm{ES}_{\mathrm{plane}}}{Q},$ (20)
$n_{\mathrm{oos}}=\overline{\mathrm{SL}}_{\mathrm{wait}}\mu_{\mathrm{oos}}N_{t}.$ (21)
2.3 Parking Spares
Parking spares respond to orders from in-plane spares, which are propagated from the failures of operational satellites. Empirically, if $N_{\mathrm{plane}}$ is large enough, a Poisson process can approximate the demand process in parking spares [29, 28]. Consequently, the demand rate for parking spares ($\lambda_{\mathrm{parking}}$) and the PMF of demand given time interval $\tau$ ($P_{\mathrm{plane}}$) are as follows:
$\lambda_{\mathrm{parking}}=\frac{\lambda_{\mathrm{plane}}\gamma_{0}N_{\mathrm{plane}}}{QN_{\mathrm{parking}}},$ (22)
$P_{\mathrm{parking}}\left(\delta;\tau\right)=\frac{\left(\lambda_{\mathrm{parking}}\tau\right)^{\delta}}{\delta!}e^{-\lambda_{\mathrm{parking}}\tau},\quad\delta\in\mathbb{Z}_{+}.$ (23)
Similar to in-plane spares, parking spares are managed under an $(s,Q)$ policy with a reorder point $s=k_{s}$ and an order quantity $Q=k_{Q}$. Orders are placed in batches of size $k_{Q}$ and fulfilled from ground spares via launch vehicles. The lead time consists of a fixed processing time, $t_{\mathrm{lau}}$, and a stochastic delay due to launch scheduling, modeled as an exponential distribution with rate parameter $\mu_{\mathrm{lau}}$ [4]. As a result, the PDF of lead time is as follows:
$f_{\mathrm{parking}}\left(\tau\right)=\mu_{\mathrm{lau}}e^{-\left(\tau-t_{\mathrm{lau}}\right)\mu_{\mathrm{lau}}}.$ (24)
Following these demand and supply relationships, the key inventory metrics for parking spares, including mean stock level ($\overline{\mathrm{SL}}_{\mathrm{parking}}$), order frequency ($n_{\mathrm{parking}}$), expected shortage ($\mathrm{ES}_{\mathrm{parking}}$), and order fill rate ($\rho_{\mathrm{parking}}$), are computed as follows [4, 5, 6]:
$\overline{\mathrm{SL}}_{\mathrm{parking}}=\int_{\tau\geq t_{\mathrm{lau}}}\left(k_{s}-\lambda_{\mathrm{parking}}\tau+\frac{k_{Q}}{2}+\frac{1}{2}\right)f_{\mathrm{parking}}\left(\tau\right)\,d\tau,$ (25)
$n_{\mathrm{parking}}=\frac{\lambda_{\mathrm{parking}}}{k_{Q}}N_{t},$ (26)
$\mathrm{ES}_{\mathrm{parking}}=\int_{\tau\geq t_{\mathrm{lau}}}\sum_{\delta\geq k_{s}}\left(\delta-k_{s}\right)P_{\mathrm{parking}}\left(\delta;\tau\right)f_{\mathrm{parking}}\left(\tau\right)\,d\tau,$ (27)
$\rho_{\mathrm{parking}}=1-\frac{\mathrm{ES}_{\mathrm{parking}}}{k_{Q}}.$ (28)
2.4 Annual Maintenance Cost and Mean Time to Disposal
Based on the inventory parameters for in‐plane and parking spares, we model the annual maintenance cost (AMC) for maintaining the satellite constellation. The AMC consists of the annual launch cost ($C_{\mathrm{lau}}$), maneuvering cost ($C_{\mathrm{maneuv}}$), manufacturing cost ($C_{\mathrm{manufac}}$), holding cost ($C_{\mathrm{hold}}$)—which together represent the cost elements for managing spare satellites by previous works [4, 5]—and the servicing cost ($C_{\mathrm{oos}}$):
$\mathrm{AMC}=C_{\mathrm{lau}}+C_{\mathrm{maneuv}}+C_{\mathrm{manufac}}+C_{\mathrm{hold}}+C_{\mathrm{oos}},$ (29)
where the individual cost components are as follows:
$C_{\mathrm{lau}}=c_{\mathrm{lau}}n_{\mathrm{parking}}N_{\mathrm{parking}},$ (30)
$C_{\mathrm{manufac}}=M_{\mathrm{fuel}}\epsilon_{\mathrm{fuel}}n_{\mathrm{plane}}N_{\mathrm{plane}}Q,$ (31)
$C_{\mathrm{manufac}}=C_{\mathrm{sat}}\lambda_{\mathrm{plane}}\gamma_{0}N_{\mathrm{plane}}N_{t},$ (32)
$C_{\mathrm{hold}}=c_{\mathrm{hold}}\left(\overline{\mathrm{SL}}_{\mathrm{parking}}QN_{\mathrm{parking}}+\overline{\mathrm{SL}}_{\mathrm{plane}}N_{\mathrm{plane}}+\overline{\mathrm{SL}}_{\mathrm{wait}}N_{\mathrm{plane}}\right),$ (33)
$C_{\mathrm{oos}}=p_{\mathrm{oos}}n_{\mathrm{oos}}N_{\mathrm{plane}}.$ (34)
Here, the annual launch cost is determined by the launch cost ($c_{\mathrm{lau}}$) and the order frequency of parking spares ($n_{\mathrm{parking}}$); the annual maneuvering cost depends on the fuel mass required for orbit raising ($M_{\mathrm{fuel}}$), the specific cost of fuel ($\epsilon_{\mathrm{fuel}}$), and the order frequency of in-plane spares ($n_{\mathrm{plane}}$); the annual manufacturing cost is governed by the manufacturing cost per satellite ($c_{\mathrm{manufac}}$), the failure rate ($\lambda_{\mathrm{plane}}$), and the fraction of in-plane spare demand met by newly launched satellites ($\gamma_{0}$); the annual holding cost depends on the holding cost rate ($c_{\mathrm{hold}}$) and the mean stock levels at parking spares ($\overline{\mathrm{SL}}_{\mathrm{parking}}$), in-plane spares ($\overline{\mathrm{SL}}_{\mathrm{plane}}$), and failed satellites awaiting service ($\overline{\mathrm{SL}}_{\mathrm{wait}}$); finally, the annual servicing cost is determined by the OOS price ($p_{\mathrm{oos}}$) and the service frequency ($n_{\mathrm{oos}}$).
Other than cost elements, the mean time from deployment to disposal of a satellite serviced $N_{\mathrm{oos}}$ times, $T_{d}$, is tracked. Suppose a satellite is disposed of after reaching the maximum number of allowable services. Then, it undergoes the following sequence: it remains as a parking spare after launch, transfers from a parking orbit to a target orbit, stays as an in-plane spare, operates, and is serviced after failure, repeatedly, until the service limit is reached.
Accordingly, $T_{d}$ can be modeled as the sum of the mean times spent in various phases: as a parking spare ($T_{\mathrm{parking}}$), during orbital transfer from parking to target orbit including RAAN alignment and orbit raising ($T_{\mathrm{trans}}$), as an in-plane spare ($T_{\mathrm{plane}}$), during operation ($T_{\mathrm{opr}}$), and while waiting for on-orbit servicing ($T_{\mathrm{oos}}$):
$T_{d}=T_{\mathrm{parking}}+T_{\mathrm{trans}}+T_{\mathrm{plane}}+T_{\mathrm{opr}}+T_{\mathrm{oos}},$ (35)
where
$T_{\mathrm{parking}}=\frac{\overline{\mathrm{SL}}_{\mathrm{parking}}}{\lambda_{\mathrm{parking}}},$ (36)
$T_{\mathrm{trans}}=\int_{\tau\geq t_{\mathrm{trans}}}\tau f_{\mathrm{plane}}\left(\tau\right)\,d\tau,$ (37)
$T_{\mathrm{plane}}=\frac{\overline{\mathrm{SL}}_{\mathrm{plane}}\left(N_{\mathrm{oos}}+1\right)}{\lambda_{\mathrm{plane}}},$ (38)
$T_{\mathrm{opr}}=\frac{\left(N_{\mathrm{oos}}+1\right)N_{t}}{\lambda_{\mathrm{sat}}},$ (39)
$T_{\mathrm{oos}}=\frac{N_{\mathrm{oos}}}{\mu_{\mathrm{oos}}}.$ (40)
2.5 Model Validation
To evaluate the accuracy of our inventory management model, we compared its parameter estimates to the mean of 100 Monte Carlo simulation results for 25 randomly selected instances. Each simulation spans 60 years and is discretized into one-day intervals. The fixed parameters and trade space for randomly sampled parameters are in Tables 2 and 3. We repeatedly sampled the varying parameters until we obtained 25 instances satisfying the following conditions:
$\rho_{\mathrm{plane}}\geq 0.98,$ (41)
$\rho_{\mathrm{parking}}\geq 0.98,$ (42)
$T_{d}/N_{t}\leq T_{\mathrm{life}},$ (43)
$s\leq Q,$ (44)
$k_{s}\leq k_{Q}.$ (45)
Here, Eqs. (41) and (42) represent the rare shortage assumptions in Section 2.1; Eq. (43) imposes the lifespan condition; and Eqs. (44) and (45) maintain the inventory’s cyclic behavior.
Table 2: Fixed parameters for validation study
Parameter, Unit Notation Value
Fuel mass conversion coefficient, M\$/kg $\epsilon_{\mathrm{fuel}}$ 0.01
Satellite production cost, M\$/satellite $c_{\mathrm{sat}}$ 0.5
Spare/failed satellite holding cost, M\$/satellite/year $c_{\mathrm{hold}}$ 0.5
Launch cost, M\$/launch $c_{\mathrm{lau}}$ 67
Specific impulse of satellite’s propulsion system, sec $I_{\mathrm{sp}}$ 1200
Propellant mass rate, kg/s $\dot{M}_{\mathrm{fuel}}$ 1.3E-6
Satellite’s dry mass, kg $M_{\mathrm{dry}}$ 150
OOS price, M\$ $p_{\mathrm{oos}}$ 2
Lifespan of satellite, years $T_{\mathrm{life}}$ 30
Table 3: Sampled trade space of varying parameters for validation study
Parameter, Unit Notation Range
Launch order processing time, weeks $t_{\mathrm{lau}}$ $\left[4,16\right]$
Mean waiting time to launch, weeks $\mu_{\mathrm{lau}}^{-1}$ $\left[4,16\right]$
Altitude of orbital planes in the constellation, km $h_{\mathrm{plane}}$ $\left[500,2000\right]$
Altitude of parking orbits, km $h_{\mathrm{parking}}$ $\left[400,1000\right]$
Inclination, deg $i$ $\left[40,80\right]$
Satellite failure rate, failures/year $\lambda_{\mathrm{sat}}$ $\left[0.01,0.2\right]$
Number of orbital planes in the constellation, planes $h_{\mathrm{plane}}$ $\left[20,40\right]$
Number of parking orbits, planes $N_{\mathrm{parking}}$ $\left[1,20\right]$
Number of satellites per orbital plane in the constellation, satellites/plane $N_{\mathrm{sat}}$ $\left[20,60\right]$
reorder point of in-plane spare, satellites $s$ $\left[1,20\right]$
Order quantity at in-plane spare, satellites $Q$ $\left[1,40\right]$
reorder point of parking spare, - $k_{s}$ $\left[1,20\right]$
Order quantity at parking spare, - $k_{Q}$ $\left[1,40\right]$
MTTR of OOS, weeks $\mu_{\mathrm{oos}}^{-1}$ $\left[1,12\right]$
Serviceable failure fraction, - $r_{\mathrm{oos}}$ $\left[0.1,0.9\right]$
Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
The absolute error evaluates rational parameters (e.g., $\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$):
$\left|\mathrm{value}_{\mathrm{sim}}-\mathrm{value}_{\mathrm{model}}\right|,$ (46)
where $\mathrm{value}_{\mathrm{sim}}$ and $\mathrm{value}_{\mathrm{model}}$ are the parameter values obtained from the simulation and the model, respectively. The relative error assesses the other parameters:
$\frac{\left|\mathrm{value}_{\mathrm{sim}}-\mathrm{value}_{\mathrm{model}}\right|}{\mathrm{value}_{\mathrm{sim}}}\times 100.$ (47)
The results in Table 4 indicate that relative errors remain below 2% and absolute errors below 1%p, meaning that the constructed model accurately estimates the parameters.
Table 4: Averaged errors of result parameters
Result Notation Mean Error
Relative Errors (%)
Mean stock level at in-plane spare $\overline{\mathrm{SL}}_{\mathrm{plane}}$ 1.95
Mean stock level at parking spare $\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.74
Mean stock level of failed satellites waiting for service $\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.85
Order frequency at in-plane spare $n_{\mathrm{plane}}$ 0.49
Order frequency at parking spare $n_{\mathrm{parking}}$ 1.88
Service frequency $n_{\mathrm{oos}}$ 0.66
Mean time to disposal of satellites serviced up to the maximum $T_{d}$ 0.35
Annual maintenance cost $\mathrm{AMC}$ 0.71
Absolute Errors (%p)
Order fill rate at in-plane spare $\rho_{\mathrm{plane}}$ 0.04
Order fill rate at parking spare $\rho_{\mathrm{parking}}$ 0.05
3 Decision-Making Scenarios and Optimization Problem Formulations
Based on the model, we develop decision-making scenarios involving two stakeholders—the constellation operator and the OOS provider—and discuss potential tension between them. This section presents the details of the decision-making contexts under consideration and their mathematical formulations as optimization problems.
3.1 Constellation Operator
The goal of the constellation operator is to maintain their system at an appropriate service level while minimizing relevant costs. Such a decision-making context can be formulated as an optimization problem, $\mathbf{P}_{1}$.
($\mathbf{P}_{1}$) Minimizing the Annual Maintenance Cost of the Constellation
$\min\ \mathrm{AMC}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right),$ (48)
subject to
$\rho_{\mathrm{plane}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{plane}}^{\mathrm{req}}$ (49)
$\rho_{\mathrm{parking}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{parking}}^{\mathrm{req}}$ (50)
$Qk_{Q}\leq Q_{\max},$ (51)
Eqs. (43)-(45),
$\mathbf{x}_{\mathrm{con}}=\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}.$ (52)
Here, the vector of decision variables, $\mathbf{x}_{\mathrm{con}}$ in Eq. (52), includes parameters related to the replenishment policy and the configuration of parking orbits, while $\mathbf{q}_{1}$ contains the remaining parameters. Equation (48) defines the objective of the problem (minimizing the annual system maintenance cost). Equations (49) and (50) ensure that the maintenance system maintains an appropriate service level. Equation (51) constrains the number of loaded satellites so that it does not exceed the launch capacity ($Q_{\max}$). Finally, $\mathcal{H}_{\mathrm{parking}}$ denotes the set of available parking orbit altitudes.
3.2 OOS Provider
The main goal of an OOS provider is to maximize profit. The annual profits (AP) of the OOS provider are defined as
$\mathrm{AP}=\left(p_{\mathrm{oos}}-c_{\mathrm{oos}}\right)n_{\mathrm{oos}}N_{\mathrm{plane}},$ (53)
where $p_{\mathrm{oos}}$ and $c_{\mathrm{oos}}$ are the unit price and cost of an OOS service, and $n_{\mathrm{oos}}N_{\mathrm{plane}}$ corresponds to the total annual number of services performed.
Given the OOS technologies, service types, and target subsystems of the spacecraft to be serviced, the OOS provider’s operational effort determines the service’s responsiveness and cost. For example, if the OOS provider allocates more resources (e.g., additional servicing spacecraft, depots, and propellant for faster orbital transfers), the servicing cost increases, but so does responsiveness. While a detailed investigation of operational rule-setting and cost estimation should be a separate research topic, here we introduce a first-cut model representing the relationship between cost and responsiveness as follows:
$c_{\mathrm{oos}}=c_{\mathrm{oos},\min}+\frac{\alpha_{1}}{\left(\mu_{\mathrm{oos}}^{-1}-\mu_{\mathrm{oos},\mathrm{ideal}}^{-1}\right)^{\alpha_{2}}}.$ (54)
Here, $c_{\mathrm{oos},\min}$ (>0>0) denotes the lower bound of the cost, $\mu_{\mathrm{oos},\mathrm{ideal}}$ (>0>0) represents the upper bound of responsiveness, and $\alpha_{1}$ (>0>0) and $\alpha_{2}$ (>0>0) are shape parameters. This function is strictly decreasing with respect to MTTR, exhibiting knee points and two asymptotes parallel to the cost and MTTR axes. Figure 3 illustrates the overall trend and the effect of each parameter.
Such a relationship, where increased effort yields greater outcomes but with a diminishing rate of increment, reflects the trade-off between service cost and responsiveness. This behavior is consistent with empirical observations in terrestrial logistics [30] and prior studies on OOS [32, 31, 33].
Figure 3: Visualization of the adopted cost-responsiveness function with the effect of each parameter (1: $c_{\mathrm{oos},\min}$, 2: $\mu_{\mathrm{oos},\mathrm{ideal}}$, 3: $\alpha_{1}$, 4: $\alpha_{2}$).
With the aid of the cost-responsiveness function, we formulate the optimization problem $\mathbf{P}_{2}$, representing the decision-making of the OOS provider:
($\mathbf{P}_{2}$) Maximizing Profits of OOS Operations
$\max\ \mathrm{AP}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right),$ (55)
subject to
$c_{oos}\left(\mu_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq p_{\mathrm{oos}},$ (56)
$\mathrm{AMC}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (57)
$\mathbf{x}_{\mathrm{oos}}=\left[p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]\in\mathbb{R}_{++}^{2}.$ (58)
Here, the vector of decision variables $\mathbf{x}_{\mathrm{oos}}$ in Eq. (58) consists of two components: the price of the OOS per service, $p_{\mathrm{oos}}$, and the service responsiveness, $\mu_{\mathrm{oos}}$. Note that $\mu_{\mathrm{oos}}$ is the inverse of the MTTR. The vector $\mathbf{q}_{2}$ contains the remaining parameters, including those associated with the maintenance strategy of the constellation operator, which served as decision variables in $\mathbf{P}_{1}$. The objective function in Eq.(55) seeks to maximize the profits. Equation (56) ensures that the OOS price is no less than the corresponding service cost. Equation (56) ensures that the OOS price is no less than the corresponding service cost. Equation (57) constrains the OOS provider to set its service charge and responsiveness within a range that prevents the constellation operator from abandoning the OOS-integrated strategy. The term $\mathrm{AMC}{\mathrm{ref}}$ serves as a reference maintenance cost for adopting OOS; it can be interpreted as the maintenance cost without OOS or as a value that includes a premium over the reference cost from an alternative, non-OOS strategy. This premium reflects the additional complexity and uncertainty introduced by OOS integration, for which the operator may require greater expected benefits.
3.3 Tension Between Constellation Operator & OOS Provider
Previous subsections introduced the optimization problems reflecting the decision-making contexts of two main stakeholders: the constellation operator and the OOS provider. Note that these two problems are highly interdependent. On one hand, if the OOS provider raises the OOS price to increase profitability, the constellation operator would rely less on OOS for system maintenance and instead deploy new satellites. On the other hand, if the constellation operator relies more on OOS, the profitability of the OOS provider would improve without changing its decisions. Balancing such tensions and designing appropriate incentives to achieve mutually beneficial outcomes that stimulate the overall space industry are crucial challenges. Indeed, the interdependence between satellite operators and OOS providers is sometimes regarded as a chicken-and-egg problem [34].
The following bi-objective optimization problem expresses the tension between the constellation operator and the OOS provider by combining $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$:
($\mathbf{P}_{3}$) Bi-Objective Optimization for Multiple Stakeholders
$\min\ \mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (59)
$\max\ \mathrm{AP}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (60)
subject to
$Qk_{Q}\leq Q_{\max},$ (61)
$c_{\mathrm{oos}}\left(\mu_{\mathrm{oos}};\mathbf{q}_{3}\right)\leq p_{\mathrm{oos}},$ (62)
$\mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (63)
Eqs. (43)-(45),
$\mathbf{x}=\left[\mathbf{x}_{\mathrm{con}},\mathbf{x}_{\mathrm{oos}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}\times\mathbb{R}_{++}^{2}.$ (64)
Here, the vector of decision variables, $\mathbf{x}$ in Eq. (64), is the concatenation of $\mathbf{x}_{\mathrm{con}}$ and $\mathbf{x}_{\mathrm{oos}}$ from $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, respectively. Solving this bi-objective optimization problem yields a family of non-dominated solutions—solutions that are not inferior to any other solution with respect to all objective functions [35]—known as the Pareto front [36, 37]. To obtain the Pareto front, we employ the non-dominated sorting genetic algorithm II (NSGA-II), a metaheuristic based on the genetic algorithm [38].
4 Case Study
This section presents a case study of the proposed bi-objective optimization problem ($\mathbf{P}_{3}$). The case study demonstrates the applicability of the proposed framework using a realistically sized constellation and representative satellite specifications. We first establish and analyze a baseline scenario, then extend the analysis through a parametric study by varying the OOS-related parameters.
4.1 Baseline Scenario
We established the baseline scenario using the constellation’s system parameters, launch service parameters, and OOS parameters, which are summarized in Tables 5, 6, and 7, respectively. The reference annual maintenance cost, $\mathrm{AMC}_{\mathrm{ref}}$, was assessed using the strategy proposed by Jakob et al. [4], which has the same supply chain as the one proposed here but without incorporating OOS, and uses the same constellation’s system and launch service parameters. The resulting optimal solution is presented in Table 8. The optimal maintenance cost is \$925.1M satisfying the service level over requirements. In addition to the scenario parameters, the decision variables and their range are summarized in Table 9.
Table 5: Constellation’s system parameters for baseline scenario
Parameter, Unit Notation Value
Inclination, deg $i$ 60
Satellite failure rate, failures/year $\lambda_{\mathrm{sat}}$ 0.2
Number of orbital planes, planes $N_{\mathrm{plane}}$ 40
Number of satellites per orbital plane, satellites/plane $N_{\mathrm{sat}}$ 40
Altitude of orbital planes, km $h_{\mathrm{plane}}$ 1200
Fuel mass conversion coefficient, \$M/kg $\epsilon_{\mathrm{fuel}}$ 0.01
Annual satellite holding cost, \$M/satellite/year $c_{\mathrm{hold}}$ 0.5
Specific impulse of satellite’s propulsion system, sec $I_{\mathrm{sp}}$ 1200
Dry mass, kg $M_{\mathrm{dry}}$ 150
Propellant mass rate, kg/s $\dot{M}_{\mathrm{fuel}}$ 1.3E-6
Production cost, \$M/satellite $c_{\mathrm{sat}}$ 0.5
Lifespace of spacecraft, years $T_{\mathrm{life}}$ 30
Service level requirement, - $\rho_{\mathrm{plane}}^{\mathrm{req}}$ 0.98
$\rho_{\mathrm{parking}}^{\mathrm{req}}$ 0.98
Annual maintenance cost threshold for adopting OOS, \$M/year $\mathrm{AMC}_{\mathrm{ref}}$ 925.1
Table 6: Launch service parameters for baseline scenario
Parameter, Unit Notation Value
Launch cost, \$M $c_{\mathrm{lau}}$ 67
Launch capacity, satellites $Q_{\max}$ 40
Launch order processing time, weeks $t_{\mathrm{lau}}$ 12
Mean waiting time to launch, weeks $\mu_{\mathrm{lau}}^{-1}$ 8
Table 7: OOS parameters for baseline scenario
Parameter, Unit Notation Value
Serviceable failures ratio, - $r_{\mathrm{oos}}$ 0.25
The upper bound of responsiveness parameter, week-1 $\mu_{\mathrm{oos},\mathrm{ideal}}$ 0.5
The lower bound of service cost, \$M $c_{\mathrm{oos},\min}$ 0.5
Shape parameters for cost-reponsiveness function, - $\alpha_{1}$ 1
$\alpha_{2}$ 1
Table 8: Optimal solution and result parameters of the baseline scenario without OOS
Result Value
Decision Variables
$s$, satellites 4
$Q$, satellites 4
$k_{s}$, - 10
$k_{Q}$, - 10
$N_{\mathrm{parking}}$, planes 6
$h_{\mathrm{parking}}$, km 795.4
Performance Outputs
$\mathrm{AMC}$, \$M/year 925.1
$C_{\mathrm{lau}}$, \$M/year 536.0
$C_{\mathrm{hold}}$, \$M/year 220.8
$C_{\mathrm{manufac}}$, \$M/year 160.0
$\rho_{\mathrm{plane}}$, % 98.0
$\rho_{\mathrm{parking}}$ % 98.3
Table 9: Decision variables and their ranges for baseline scenario
Decision Variable, Unit Notation Range
$\mathbf{x}_{\mathrm{con}}$
Reorder point of in-plane spares, satellites $s$ $\left[1,20\right]$
Order quantity of in-plane spares, satellites $Q$ $\left[1,Q_{\max}\right]$
Reorder point of parking spares, - $k_{s}$ $\left[1,20\right]$
Order quantity of parking spares, - $k_{Q}$ $\left[1,Q_{\max}\right]$
Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
Number of parking orbits, planes $N_{\mathrm{parking}}$ $\left[1,20\right]$
Altitude of parking orbits, km $h_{\mathrm{parking}}$ $\left[500,1000\right]$
$\mathbf{x}_{\mathrm{oos}}$
OOS price, \$M $p_{\mathrm{oos}}$ $\left[c_{\mathrm{oos},\min},c_{\mathrm{oos},\min}+5\right]$
OOS MTTR, weeks $\mu_{\mathrm{oos}}^{-1}$ $\left(\mu_{\mathrm{oos},{\mathrm{ideal}}}^{-1},12\right]$
Figure 4: Solutions of the baseline scenario scattered in the objective space (color: OOS price).
Table 10: Decision variables and resultant parameters of the Pareto-optimal solutions: baseline scenario
Result, Unit Value or Range
Decision Variables
$\mathbf{x}_{\mathrm{con}}$
$s$, satellites 3
$Q$, satellites 4
$k_{s}$, - 6
$k_{Q}$, - 10
$N_{\mathrm{parking}}$, planes 7
$h_{\mathrm{parking}}$, km [674.3, 700.4]
$N_{\mathrm{oos}}$, - 4
$\mathbf{x}_{\mathrm{oos}}$
$p_{\mathrm{oos}}$, \$M [0.6, 2.3]
$\mu_{\mathrm{oos}}^{-1}$, weeks [10.6, 12.0]
Performance Outputs
$\mathrm{AMC}$, \$M/year [790.6, 925.0]
$\mathrm{AP}$, \$M/year [0, 134.5]
$C_{\mathrm{lau}}$, \$M/year 402.4
$C_{\mathrm{hold}}$, \$M/year [211.4, 212.6]
$C_{\mathrm{manufac}}$, \$M/year 120.1
$C_{\mathrm{oos}}$, \$M/year [48.1, 182.4]
$T_{d}/N_{t}$, years [29.2, 29.3]
$\rho_{\mathrm{plane}}$, % [98.0, 98.1]
$\rho_{\mathrm{parking}}$, % 98.1
$\gamma_{0}$, % 75.1
The obtained optimal solutions are scattered in the objective space in Fig. 4. The $x$-axis and $y$-axis in Fig. 4 represent the first objective (i.e., the constellation operator’s annual maintenance cost, $\mathrm{AMC}$) and the second objective (i.e., the OOS provider’s annual profit $\mathrm{AP}$), respectively. Since the first objective is minimized and the second is maximized, the utopia point is located at the upper-left corner of the figure, marked by a red star.
The values and ranges of the decision variables, along with the resulting parameters of the optimal solutions, are summarized in Table 10. Interestingly, all solutions share identical or very similar values for the individual components of $\mathbf{x}_{\mathrm{con}}$, decision variables for the constellation operator. Notably, the optimal value of $N_{\mathrm{oos}}$ is 4, its upper bound, suggesting that once OOS is adopted, performing as many OOS operations as possible is preferable to alternative maintenance policies in this bi-objective optimization framework with the given parameters. Consequently, $T_{d}$ approaches the value of $T_{\mathrm{life}}$, 30 years.
In contrast, the OOS provider’s decision variables vary significantly, especially $p_{\mathrm{oos}}$, determining the position of each solution in the objective space. Specifically, the decisions are significantly influenced by the OOS price: as the price increases, the constellation operator’s expenditure rises, while the OOS provider’s profitability improves. On the other hand, the MTTR value is at its worst-case (upper-bound) value of 12, minimizing the OOS provider’s cost associated with service operations.
Compared with the optimal without-OOS strategy presented in Table 8, the launch cost is reduced from \$536M to \$402.4M (approximately 25% reduction), and the manufacturing cost decreases from \$160M to \$120.1M (also about 25% reduction). These reductions result from leveraging OOS, represented by $\gamma_{0}=75\%$, indicating that only 75% of failures are handled by newly launched satellites. These values remain consistent across all Pareto-optimal solutions since the constellation operator’s maintenance strategy remains unchanged, and the responsiveness of the OOS service varies only within a narrow range. Thus, the annual OOS cost, ranging from \$48.1M to \$182.4M, determines the cost-effectiveness of the OOS-integrated strategy. The Pareto-optimal solutions form a line with a slope of 1, as both objectives are linear with respect to $p_{\mathrm{oos}}$ and share the same proportional coefficient, $n_{\mathrm{oos}}N_{\mathrm{oos}}$—the total annual number of OOS service calls.
Overall, once the OOS technologies, service types, and target spacecraft subsystems are defined, the balance between the constellation operator’s expenditure and the OOS provider’s profitability across the non-dominated solutions depends primarily on the service charge. In contrast, the constellation operator’s optimal maintenance strategy remains unchanged under varying prices.
4.2 Parametric Study
[…]
Table 11: Cost-responsiveness function parameters for parametric study instances
$r_{\mathrm{oos}}$ $c_{\mathrm{oos},\min}$ $\mu_{\mathrm{oos},\mathrm{ideal}}$ $\alpha_{1}$ $\alpha_{2}$
Instance ID (-) (\$M) (week-1) (-) (-)
Instance 0 (Baseline) 0.25 0.5 0.5 1 1
Instance 1-1 0.5 0.5 0.5 1 1
Instance 1-1 0.1 0.5 0.5 1 1
Instance 2-1 0.25 1 0.5 1 1
Instance 2-2 0.25 0.25 0.5 1 1
Instance 3-1 0.25 0.5 1 1 1
Instance 3-2 0.25 0.5 0.25 1 1
Instance 4-1 0.25 0.5 0.5 2 1
Instance 4-2 0.25 0.5 0.5 0.5 1
Instance 5-1 0.25 0.5 0.5 1 2
Instance 5-2 0.25 0.5 0.5 1 0.5
4.2.1 Serviceable Failures Ratio: $r_{\mathrm{oos}}$
We first investigate the effect of the serviceable-failure ratio ($r_{\mathrm{oos}}$). In Instances 1-1 and 1-2, the $r_{\mathrm{oos}}$ values are set as 0.5 and 0.1, respectively. The results are shown in Fig. 5(a) and Table 12.
(a) Instance 1-X: variations in $r_{\mathrm{oos}}$.
(b) Instance 2-X: variations in $c_{\mathrm{oos},\min}$.
(c) Instance 3-X: variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$.
(d) Instance 4-X: variations in $\alpha_{1}$.
(e) Instance 5-X: variations in $\alpha_{2}$.
Figure 5: Pareto fronts of each Instance in the objective space.
Table 12: Parametric study results: variations in $r_{\mathrm{oos}}$
Result, Unit Instance 1-1 Instance 1-2
($r_{\mathrm{oos}}=0.5$) ($r_{\mathrm{oos}}=0.1$)
Decision Variables
$\mathbf{x}_{\mathrm{con}}$
$s$, satellites 3 4
$Q$, satellites 3 4
$k_{s}$, - 5 9
$k_{Q}$, - 13 10
$N_{\mathrm{parking}}$, planes 7 6
$h_{\mathrm{parking}}$, km [638.3, 700.4] [818.3, 823.3]
$N_{\mathrm{oos}}$, - 4 [3, 4]
$\mathbf{x}_{\mathrm{oos}}$
$p_{\mathrm{oos}}$, \$M [0.6, 2.3] [0.6, 2.2]
$\mu_{\mathrm{oos}}^{-1}$, weeks [10.4, 12.0] [10.4, 12.0]
Performance Outputs
$\mathrm{AMC}$, \$M/year [661.9, 925.0] [873.4, 925.0]
$\mathrm{AP}$, \$M/year [0, 263.1] [0, 51.7]
$C_{\mathrm{lau}}$, \$M/year 283.7 482.4
$C_{\mathrm{hold}}$, \$M/year [194.9, 197.4] [220.2, 220.9]
$C_{\mathrm{manufac}}$, \$M/year 82.6 144.0
$C_{\mathrm{oos}}$, \$M/year [93.0, 356.3] [19.4, 71.0]
$T_{d}/N_{t}$, years [29.2, 29.4] [23.5, 29.4]
$\rho_{\mathrm{plane}}$, % [98.0, 98.1] 98.0
$\rho_{\mathrm{parking}}$, % 98.0 98.4
$\gamma_{0}$, % 51.6 90.0
From Fig. 5(a), we observed that increasing and decreasing $r_{\mathrm{oos}}$ shift the Pareto-optimal solutions toward the upper-left and lower-right directions, respectively—corresponding to movements toward and away from the utopia point. Moreover, the Pareto fronts of all instances appear nearly parallel. This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance, although differences among instances exist, as shown in Table 12.
Instance 1-1 has $N_{\mathrm{oos}}$ fixed at its upper bound, whereas Instance 1-2 has $N_{\mathrm{oos}}$ values of 3 and 4, slightly detached from the upper bound. However, despite the variation in $N_{\mathrm{oos}}$, the $\gamma_{0}$ value of Instance 1-2 remains at 0.9. This is because $r_{\mathrm{oos}}=0.1$ is sufficiently small that changes in $N_{\mathrm{oos}}$ between 3 and 4 have only a minimal effect on $\gamma_{0}$ (Eq. (9)).
Interestingly, although the OOS provider’s allowable decision ranges—those of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$—remain almost the same as in the baseline scenario, the ranges of the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, differ significantly. This implies that the constellation operator must reconsider its strategy carefully when $r_{\mathrm{oos}}$ changes substantially to maintain non-dominated solutions, even though the allowable responsiveness and OOS charge remain nearly unchanged.
By examining the cost components and $\gamma_{0}$, we found that the deployment of new satellites decreases as $r_{\mathrm{oos}}$ decreases. Specifically, from Instance 1-1 to Instances 0 and 1-2, the values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$ increase from \$283.7M to \$482.4M, from \$82.6M to \$144.0M, and from 51.6% to 90.0%, respectively, a natural outcome of the reduced contribution of OOS services.
4.2.2 Shift Parameters of Cost-Responsiveness Function: $c_{\mathrm{oos},\min}$ & $\mu_{\mathrm{oos},\mathrm{ideal}}$
Next, we conducted experiments by varying the shift parameters of the cost-responsiveness function in Eq. (54), $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$. The resulting solutions are illustrated in Figs. 5(b) & 5(c), and the detailed values and ranges of the decision variables and result parameters are provided in Table 13.
Table 13: Parametric study results: variations in $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$
Result, Unit Instance 2-1 Instance 2-2 Instance 3-1 Instance 3-2
($c_{\mathrm{oos},\min}=1$) ($c_{\mathrm{oos},\min}=0.25$) ($\mu_{\mathrm{oos},\mathrm{ideal}}=1$) ($\mu_{\mathrm{oos},\mathrm{ideal}}=0.25$)
Decision Variables
$\mathbf{x}_{\mathrm{con}}$
$s$, satellites 3
$Q$, satellites 4
$k_{s}$, - 6
$k_{Q}$, - 10
$N_{\mathrm{parking}}$, planes 7
$h_{\mathrm{parking}}$, km [685.3, 700.4] [675.5, 700.4] [676.5, 700.4] [656.5, 700.4]
$N_{\mathrm{oos}}$, - 4
$\mathbf{x}_{\mathrm{oos}}$
$p_{\mathrm{oos}}$, \$M [1.1, 2.3] [0.4, 2.3] [0.6, 2.3] [0.6, 2.3]
$\mu_{\mathrm{oos}}^{-1}$, weeks [10.8, 12.0] [10.3, 12.0] [9.3, 12.0] [11.2, 12.0]
Performance Outputs
$\mathrm{AMC}$, \$M/year [830.5, 925.0] [770.7, 925.0] [789.9, 925.0] [792.7, 925.0]
$\mathrm{AP}$, \$M/year [0, 94.6] [0, 154.4] [0.1, 135.2] [0.1, 132.4]
$C_{\mathrm{lau}}$, \$M/year 402.4
$C_{\mathrm{hold}}$, \$M/year [211.1, 212.6] [211.1, 212.6] [210.4, 212.4] [211.8, 212.5]
$C_{\mathrm{manufac}}$, \$M/year 120.1
$C_{\mathrm{oos}}$, \$M/year [87.8, 182.3] [28.0, 182.5] [47.8, 182.9] [50.0, 182.3]
$T_{d}/N_{t}$, years [29.2, 29.3] [29.1, 29.3] [29.2, 29.3] [29.2, 29.3]
$\rho_{\mathrm{plane}}$, % 98.0 [98.0, 98.1] [98.0, 98.1] [98.0, 98.1]
$\rho_{\mathrm{parking}}$, % 98.1
$\gamma_{0}$, % 75.1
We first observed from Fig. 5(b) that increasing and decreasing $c_{\mathrm{oos},\min}$ shift the Pareto-optimal solutions toward the lower-right and upper-left directions, respectively. Figure 5(c) shows that variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have the opposite effect, but its impact is insignificant. This is because the range of optimal $\mu_{\mathrm{oos}}$ lies in the lower-right tail of the cost-responsiveness curve shown in Fig. 3, where translation along the $x$-axis has only a limited effect on $c_{\mathrm{oos}}$.
Similar to the results of the previous instances, the Pareto fronts of all cases appear nearly parallel. This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance. In addition, we can see that the constellation operator’s decision variables remain the same as those in the baseline scenario. However, the minimum values of the optimal $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ shift in Instance 2-1 and Instance 2-2, respectively, due to the translation of the cost-responsiveness curve along the $y$- and $x$-axes.
Because the constellation operator’s decision variables remain identical to those in the baseline scenario, the dependence on OOS does not change, as reflected by the unchanged values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$. Nevertheless, variations in service charge and responsiveness lead to changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, through slight adjustments in $C_{\mathrm{hold}}$ and more substantial changes in $C_{\mathrm{oos}}$.
4.2.3 Shape Parameters of Cost-Responsiveness Function: $\alpha_{1}$ & $\alpha_{2}$
Lastly, we vary the shape parameters of the cost-responsiveness function, $\alpha_{1}$ and $\alpha_{2}$. The resulting solutions are presented in Figs. 5(d) & 5(e) and Table 14.
Table 14: Parametric study results: variations in $\alpha_{1}$ and $\alpha_{2}$
Result, Unit Instance 4-1 Instance 4-2 Instance 5-1 Instance 5-2
($\alpha_{1}=2$) ($\alpha_{1}=0.5$) ($\alpha_{2}=2$) ($\alpha_{2}=0.5$)
Decision Variables
$\mathbf{x}_{\mathrm{con}}$
$s$, satellites 3
$Q$, satellites 4
$k_{s}$, - 6
$k_{Q}$, - 10
$N_{\mathrm{parking}}$, planes 7
$h_{\mathrm{parking}}$, km [675.1, 700.4] [683.7, 700.4] [672.1, 700.4] [684.5, 700.4]
$N_{\mathrm{oos}}$, - 4
$\mathbf{x}_{\mathrm{oos}}$
$p_{\mathrm{oos}}$, \$M [0.7, 2.3] [0.6, 2.3] [0.5, 2.3] [0.8, 2.3]
$\mu_{\mathrm{oos}}^{-1}$, weeks [11.7, 12.0] [7.7, 10.9] [6.8, 9.3] [11.5, 12.0]
Result Parameters
$\mathrm{AMC}$, \$M/year [798.6, 925.0] [786.1, 925.0] [781.8, 925.0] [807.9, 925.0]
$\mathrm{AP}$, \$M/year [0.0, 126.5] [0.1, 139.0] [0.0, 143.2] [0.0, 117.2]
$C_{\mathrm{lau}}$, \$M/year 402.4
$C_{\mathrm{hold}}$, \$M/year [212.2, 212.6] [209.1, 211.6] [208.4, 210.4] [212.0, 212.6]
$C_{\mathrm{manufac}}$, \$M/year 120.1
$C_{\mathrm{oos}}$, \$M/year [55.9, 182.3] [45.1, 184.4] [42.0, 185.5] [65.2, 182.3]
$T_{d}/N_{t}$, years [29.2, 29.3] [28.9, 29.2] [28.9, 29.1] [29.2, 29.3]
$\rho_{\mathrm{plane}}$, % [98.0, 98.1] 98.0 [98.0, 98.1] [98.0, 98.1]
$\rho_{\mathrm{parking}}$, % 98.1
$\gamma_{0}$, % 75.1
From Figs. 5(d) & 5(e), we observed that an increase in $\alpha_{1}$ shifts the Pareto-optimal solutions toward the lower-right direction, whereas a change in $\alpha_{2}$ produces the opposite shift. The trend associated with $\alpha_{1}$ is intuitive, as increasing $\alpha_{1}$ raises the OOS cost for a given $\mu_{\mathrm{oos}}$. However, the effect of $\alpha_{2}$ depends on the value of $\mu_{\mathrm{oos}}$: when $\mu_{\mathrm{oos}}$ lies near the lower-right tail of the cost-responsiveness curve (i.e., for small $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ increases the OOS cost; conversely, when $\mu_{\mathrm{oos}}$ lies near the upper-left tail (i.e., for large $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ decreases the OOS cost. Because the current Pareto-optimal solutions occupy the region near the lower-right tail of the cost-responsiveness curve, variations in $\alpha_{2}$ move the overall solutions in the opposite direction of $\alpha_{1}$.
As with the previous instances, the Pareto fronts across all cases remain nearly parallel, for the same underlying reasons discussed earlier. Notably, the minimum values of $p_{\mathrm{oos}}$ vary between instances, whereas the maximum values remain almost constant, consistent with the previously observed patterns. In contrast to earlier cases, the maximum values of $\mu_{\mathrm{oos}}^{-1}$ change when either $\alpha_{1}$ decreases or $\alpha_{2}$ increases, as these adjustments make the cost-responsiveness function more favorable in shape, allowing greater responsiveness to be achieved with only a marginal increase in OOS cost.
Consistent with Instances 2-X and 3-X, the constellation operator’s decision variables are identical to those in the baseline scenario. As a result, the level of dependence on OOS remains unchanged, as reflected in the constant values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$. However, variations in the service charge and responsiveness still lead to corresponding changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$.
4.3 Results Summary and Strategic Implications
Through the case study, we observed Pareto-efficient solutions of $\mathbf{P}_{3}$ with variations in OOS-related parameters ($r_{\mathrm{oos}}$, $c_{\mathrm{oos},\min}$, $\mu_{\mathrm{oos},\mathrm{ideal}}$, $\alpha_{1}$, $\alpha_{2}$), while keeping the constellation and launch service parameters fixed. The summary of the results and their implications is provided below.
Given the constellation and launch service parameters, as well as the fraction of serviceable failures ($r_{\mathrm{oos}}$), Pareto-optimal solutions share a common maintenance strategy adopted by the constellation operator. This observation suggests that, to remain Pareto-efficient, there exists a dominant maintenance strategy for the satellite constellation. However, each Pareto-efficient solution differs in the OOS provider’s decision variables, particularly the OOS price $p_{\mathrm{oos}}$, which governs the location of the Pareto-optimal solution in the objective space. As $p_{\mathrm{oos}}$ increases, the constellation operator’s annual maintenance cost ($\mathrm{AMC}$) rises, while the OOS provider’s annual profit ($\mathrm{AP}$) also increases. Therefore, $p_{\mathrm{oos}}$ acts as a balancing parameter between the constellation operator’s expenditure and the OOS provider’s profits. Although the responsiveness of OOS ($\mu_{\mathrm{oos}}$) varies among different Pareto-optimal solutions, the variations remain within a limited range and have only minor effects on the results, since they lie within the right-bottom tail of the cost-responsiveness curve shown in Fig. 3. Consequently, the Pareto-optimal solutions form a linear trend with a slope of 1, cut by lines of $\mathrm{AP}=0$ (as the OOS cost cannot exceed the OOS price) and $\mathrm{AMC}=\mathrm{AMC}_{\mathrm{ref}}$ (due to the maintenance cost threshold set by the constellation operator). The slope of 1 is attributed to variations in $p_{\mathrm{oos}}$ affecting $\mathrm{AMC}$ linearly via Eq. (34) and $\mathrm{AP}$ linearly via Eq. (53), both with coefficients of $n_{\mathrm{oos}}N_{\mathrm{plane}}$.
Among the OOS-related parameters, the only factor that alters the constellation operator’s maintenance strategy is $r_{\mathrm{oos}}$, the serviceable failure ratio. When $r_{\mathrm{oos}}$ is sufficiently high, Pareto-optimal solutions consistently yield the maximum value of $N_{\mathrm{oos}}$, indicating that the constellation operator fully utilizes the OOS capability. However, as $r_{\mathrm{oos}}$ decreases significantly, $N_{\mathrm{oos}}$ may deviate from its upper bound, since it has little impact on the fraction of failures replaced by newly launched satellites, represented by the ratio $\gamma_{0}$ (or equivalently, the fraction of failed satellites serviced by OOS, $1-\gamma_{0}$). As a result, increasing $r_{\mathrm{oos}}$ raises the fraction of failures handled by OOS, thereby increasing the proportion of the total annual maintenance cost attributed to OOS. Conversely, when $r_{\mathrm{oos}}$ decreases, the opposite trend occurs.
Regarding the cost-responsiveness function, variations in the shift parameters—the ideal OOS cost ($c_{\mathrm{oos},\min}$) and responsiveness ($\mu_{\mathrm{oos},\mathrm{ideal}}$)—minimally affect the decision variables of the Pareto-optimal solutions. Since the Pareto-optimal solutions lie within the right-bottom tail of the cost-responsiveness curve, changes in $c_{\mathrm{oos},\min}$ directly shift the OOS cost, whereas variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have negligible influence on the overall cost and results. In contrast, variations in the shape parameters $\alpha_{1}$ and $\alpha_{2}$ significantly affect the efficient ranges of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ by reshaping the overall curvature of the cost-responsiveness curve. Increasing $\alpha_{1}$ raises the overall cost for a given responsiveness, shifting the Pareto-optimal solutions toward the degrading direction. Conversely, because the Pareto-optimal solutions are located near the right-bottom tail of the curve, increasing $\alpha_{2}$ moves the Pareto-optimal solutions toward the improving direction.
5 Conclusions
This study introduces an OOS-integrated maintenance strategy for large-scale satellite constellations, in which operators procure spare satellites simultaneously by deploying new units and restoring failed satellites using OOS. A three-echelon supply chain, consisting of in-plane spares (first), parking spares (second), and ground spares (third), is considered. The system forms a closed-loop structure connected to the in-plane spares, where failed satellites are recovered through OOS, and the restored units are used to replenish the in-plane spares.
On this foundation, we develop analytical models for inventory parameters that represent the average behavior of the supply chain over an infinite time horizon, under parametric replenishment policies. Based on these models, we introduce optimization problems relevant to two key stakeholders: the constellation operator and the OOS provider. To capture the OOS provider’s decision-making context, we further introduce a functional relationship between OOS cost and responsiveness, characterized by a strictly decreasing curve with a knee point and two asymptotes parallel to the coordinate axes, implying cost and performance bounds. Finally, we formulate each stakeholder’s problem as an individual optimization model, and then integrate them into a bi-objective optimization framework to identify the family of Pareto-optimal (ideal) solutions. Finally, the developed models and optimization framework are applied to realistic large-scale constellation maintenance scenarios under various OOS-related parameters.
There are several promising avenues for future research. First, this study constrains the cost-responsiveness relationship to a specific functional form characterized by two shift and two shape parameters. Although this form can be applied to many scenarios, the resulting insights are not universal. Moreover, since these four parameters are defined based on current estimates, and OOS technology is still emerging, the relationship between cost and performance remains uncertain and likely to evolve. Therefore, further investigation into this relationship would enhance the relevance and generality of the results presented here.
Second, we assume that a recovered system returns to an “as-good-as-new” state. In practice, however, recovered satellites are likely to experience some degree of degradation. This factor could affect the effectiveness of OOS as part of the overall maintenance strategy, making it an important topic for future study.
Third, future research may also explore more detailed and realistic decision-making contexts and relevant solution concepts. Although the current solution concept, Pareto optimality, provides valuable theoretical insights, such idealized solutions are not always directly attainable in practice. Alternative frameworks may better represent real-world interactions. For instance, in situations where the constellation operator and OOS providers act in self-interest, or where multiple OOS providers compete for service opportunities, game-theoretic approaches such as the Nash equilibrium could yield more practical and policy-relevant insights.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-02213804 & RS-2025-00563732). The authors thank the NRF for the support of this work.

### v2

@@
-On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation
+On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation00footnotetext: Presented as Paper 2026-2631 at the AIAA SciTech Forum, Orlando, FL, January 12–16, 2026
 Abstract
@@
 We develop an inventory management model with a parametric replenishment policy for the maintenance process, which can evaluate the performance of the satellite constellation system.
-Based on this model, we formulate two single-objective optimization problems representing the decision-making contexts of two main stakeholders—the constellation operator and the OOS provider—and a bi-objective optimization problem that can reflect the tension between the two.
-A case study of the OOS-supported maintenance for a real-world-scale constellation provides valuable insights that help explain the behaviors of the stakeholders.
+Based on this model, we formulate the interaction between the constellation operator—who seeks to maintain the required service level of the constellation while minimizing maintenance cost—and the OOS provider—who seeks to maximize profit by selecting service price and performance levels—as a bi-objective optimization problem and identify the corresponding Pareto-optimal solutions.
+A case study based on real-world-scale constellation and launch service shows that, relative to the benchmark strategy without OOS, the OOS-integrated solutions can reduce annual maintenance cost by up to 14.5%, while reducing annual launch and manufacturing costs by approximately 25% each and maintaining the required service levels.
+The results further show that, within a given scenario, the Pareto-optimal set is generally generated by an almost invariant maintenance strategy on the constellation operator side, whereas most variation along the Pareto frontier is driven by pricing decisions of the OOS provider side.
+Among the OOS-related parameters, the fraction of failures that can be recovered through OOS has the strongest structural effect.
 Nomenclature
@@
 $V_{\mathrm{e}}$  = exhaust velocity of satellite propulsion system, km/s
-$\mathbf{x}_{\mathrm{con}}$  = decision variables vector of constellation operator
-$\mathbf{x}_{\mathrm{oos}}$  = decision variables vector of OOS
+$\mathbf{x}$  = decision variables vector of optimization problem
 $\alpha_{1},\alpha_{2}$  = shape parameters of cost-responsiveness function, -
@@
 Satellite mega-constellations have played a pivotal role in shaping the modern space industry, driven by the demand for global coverage services and advances in space technologies.
-These driving forces have encouraged active participation from major players, including SpaceX’s Starlink, Amazon’s Project Kuiper, and Eutelsat’s OneWeb [1, 2].
+These driving forces have encouraged active participation from major players, including SpaceX’s Starlink, Amazon Leo, and Eutelsat’s OneWeb [1, 2].
 This participation has expanded the market and established mega-constellations as a critical component in the ecosystem of the space industry.
-From a system maintenance perspective, satellite mega-constellations expose operators to more frequent component failures (i.e., satellite failures), primarily due to the reduced reliability of individual satellites and the vastly increased number of units in operation.
-Addressing these challenges necessitates a systematic maintenance strategy, specifically, a well-structured logistics framework for effective failure response [3].
-Recent studies, including those by Jakob et al.
+From an operational perspective, mega-constellations in low Earth orbit (LEO) pose challenges throughout the lifecycle that differ substantially from those of traditional constellations in geostationary equatorial orbit (GEO).
+The large number of satellites, together with the relatively small size of individual spacecraft and tight constraints on power, performance, and maneuvering capability, complicates system operation and management.
+Prior research has addressed multiple lifecycle phases, including pattern design [3, 4, 5, 6, 7, 8, 9, 10, 11, 13, 12, 14, 15], deployment [16, 17, 18, 19, 20, 21], mission operations [22, 23, 24, 25], and end-of-life management [26, 27].
+This study focuses on maintenance.
+In mega-constellations, component-level failures (i.e., in-orbit failures of individual satellites) may occur more frequently, not only because more satellites are in operation, but also because individual satellites are often designed under tighter mass, cost, and redundancy constraints.
+In addition, rapid technology refresh cycles may require agile integration of payload and spacecraft upgrades.
+For example, Starlink satellites are reported to have an operational lifetime of about five years, a wet mass of several hundred kilograms, and a planned constellation size of 42,000 [28].
+In contrast, current Intelsat—a GEO-based telecommunication company—satellites can exceed 6 metric tons, are expected to remain operational for more than 15 years [29], and are deployed in fleets of only a few tens of satellites.
+These differences motivate the development of a systematic maintenance strategy tailored to mega-constellations, specifically a structured logistics framework for effective failure response [30].
+Recent studies, including Jakob et al.
 [31] and Kim et al.
-[5], have proposed strategies and corresponding assessment tools to tackle this problem.
+[32], have proposed maintenance strategies and corresponding assessment tools for satellite mega-constellations.
 Specifically, Jakob et al.
@@
 [32].
-Beyond deploying spare satellites, on-orbit servicing (OOS) can recover a satellite’s operational capability and is beginning to be commercialized [8, 9, 10, 11].
-Historically, OOS has targeted heavy space systems, such as the International Space Station, the Hubble Space Telescope, and communication satellites at geosynchronous equatorial orbit, whose life extension offers significant monetary value.
-Numerous studies have quantified OOS’s benefits for single large systems [12, 13, 14, 15, 16, 17, 18], and operational successes, such as the MEV-1 mission [19], underscore its practical value.
-More recently, researchers have begun to consider OOS for low Earth orbit (LEO) constellations, where the rapid growth of space activities presents new opportunities.
-Although each individual satellite is relatively inexpensive and the added value of servicing a single unit is smaller compared to larger space systems, the overall scale of constellations continues to attract players in the space industry.
-Luu and Hastings [20] compared two strategies—deploying spare satellites and using OOS—for maintaining satellite mega-constellations.
-However, existing research has not yet examined spare satellites and OOS simultaneously as part of a unified maintenance strategy for satellite mega-constellations.
-To address this gap, we propose a novel maintenance strategy for satellite constellations that integrates OOS with the deployment of spare satellites.
-Within this strategy, failed satellites are not only replaced by spare satellites but can also be repaired through OOS if the failure is recoverable.
-To evaluate the performance of the proposed strategy at a strategic level, we construct the corresponding supply chain and develop an assessment tool based on an inventory management model.
-Building on this foundation, we introduce relevant decision-making scenarios, formulate them as optimization problems, and conduct a case study to demonstrate the applicability of the proposed approach and provide actionable insights.
-The remainder of this paper is organized into five sections.
-Section 2 introduces the modeling assumptions and constructs the supply chain and inventory management framework for evaluating the OOS-integrated maintenance strategy.
-Section 3 describes the decision-making scenarios, while Section 4 presents a case study that illustrates the application of these scenarios.
+Beyond replacing failed satellites with new spare units, on-orbit servicing (OOS) can restore operational capability and is beginning to be commercialized [34, 35, 36, 37].
+Historically, OOS has focused on high-value space systems such as the International Space Station, the Hubble Space Telescope, and communication satellites in GEO, for which life extension can provide substantial economic value.
+Numerous studies have assessed the benefits of OOS for single large systems [38, 39, 40, 41, 42, 43, 44], and operational missions such as MEV-1 [45] indicate its practical relevance.
+More recently, studies on OOS for servicing multiple satellites have examined technology and architecture alternatives [50, 46, 51, 47, 52, 48, 49], as well as the optimization of service planning and system design [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66].
+However, this body of literature primarily adopts the perspective of the OOS provider, focusing on how OOS systems—an emerging and still difficult-to-realize concept—should be designed and operated.
+By contrast, the customer-side perspective, namely how constellation operators should incorporate OOS into their system, remains relatively underexplored.
+For example, Luu and Hastings [52] compared two maintenance strategies for satellite mega-constellations from the constellation operator’s perspective: deploying spare satellites or relying on OOS.
+Nevertheless, to our knowledge, prior research has not jointly considered spare satellites and OOS within a unified maintenance strategy for satellite mega-constellations.
+To address this gap, this paper proposes a maintenance strategy for satellite constellations that integrates OOS with spare satellite management.
+Under this strategy, failed satellites may be replaced by spare units or recovered through OOS when the failure is recoverable.
+To evaluate the proposed strategy, we develop an assessment tool based on an inventory management model.
+Building on this model, we define a decision-making scenario involving two stakeholders: the constellation operator, as the customer of OOS, and the OOS provider.
+We then formulate these settings as optimization problems and conduct a case study.
+The contributions of this work are twofold.
+First, we propose an inventory control-based framework for integrating OOS into mega-constellation maintenance systems.
+The framework jointly considers spare-based replacement and recovery through OOS, and incorporates the resulting closed-loop structure into the inventory management model.
+To our knowledge, this closed-loop feature has not been addressed in prior work [31, 32].
+Second, building on this framework, we formulate an optimization problem and present a case study to demonstrate its applicability and derive managerial insights for relevant decision makers.
+In particular, we conduct a parametric analysis of OOS cost and performance, both of which remain uncertain as OOS capabilities continue to emerge.
+From the customer perspective (i.e., constellation operators), the analysis guides how operators should respond to the emergence of OOS in the market.
+From the customer perspective (i.e., constellation operators), the analysis clarifies how operators should respond to the emergence of OOS in the market.
+From the provider perspective (i.e., OOS providers), it identifies the key factors that make OOS offerings attractive to customers and examines how customers may respond to such offerings.
+The remainder of this paper is organized as follows.
+Section 2 outlines the modeling assumptions and develops the supply chain and inventory management framework to evaluate the OOS-integrated maintenance strategy.
+Section 3 validates the proposed model by comparing its results with those obtained from inventory simulations and assesses its robustness to key assumptions through analyses under relaxed assumptions.
+Section 4 presents the decision-making scenarios and corresponding optimization problem formulations, and Section 5 provides a case study illustrating their application.
 Finally, Section 6 concludes the paper.
 2 Model Construction
-This section outlines the proposed maintenance strategy and the evaluation framework.
-We begin by introducing the supply chain and the underlying assumptions.
-Next, we present the parametric models that describe the mean behavior of the inventory at each echelon of the supply chain and construct the maintenance cost and system performance parameters.
-Finally, we assess the model’s accuracy by comparing its results with those obtained from discrete-time simulations.
+This section presents the proposed maintenance strategy and evaluation framework.
+We first describe the concept of operations of the maintenance system, together with the associated supply chain and its underlying assumptions.
+We then present the parametric models to characterize the mean inventory behavior at each echelon of the supply chain and relate these models to the maintenance cost and other system performance parameters.
+2.1 Concept of Operations
+Before presenting the detailed assumptions and inventory models, we briefly describe the maintenance system. Figure 1 provides an overview of the proposed maintenance concept.
+Figure 1: Illustration of the proposed OOS-integrated maintenance strategy.
+We assume that the target constellation follows a Walker Delta pattern [67], where all orbital planes are circular, share the same inclination, and are uniformly distributed over the right ascensions of the ascending nodes (RAANs).
+Satellites within each orbital plane are also assumed to be uniformly distributed along, and all satellites in the constellation are assumed to be homogeneous.
+Although the formulation is developed for this symmetric configuration, it can be adapted with minimal effort to constellations that deviate from it, for example, by allowing nonzero eccentricity in each orbital plane or nonuniform distribution of orbital planes over the RAAN domain.
+In such cases, the model can still be used to track the mean behavior of the spare inventory, although deviations from the mean would require additional study.
+To address failures of operational satellites, we consider three echelons of spare satellites, following Jakob et al.
+[31].
+The first echelon consists of in-plane spares, which are distributed across the (near-)target orbits (i.e., the operational planes of the constellation).
+When an operational satellite fails, its slot is filled by an in-plane spare.
+The second echelon consists of parking spares, which replenish the in-plane spares.
+These parking orbits have the same inclination as the target orbits and are uniformly distributed over RAANs, but they are located at lower altitudes.
+A parking spare replenishes an in-plane spare by performing an orbit-raising maneuver once the corresponding parking and target orbits become aligned in RAAN due to the differential RAAN drift induced by the altitude difference, as discussed later.
+The third echelon consists of ground spares, which replenish the parking spares.
+Ground spares are newly manufactured satellites that are delivered to the parking orbits by launch vehicles.
+The associated replenishment flows of newly manufactured satellites are shown as blue arrows in Fig.
+1.
+In addition to the replenishment flow of newly manufactured satellites, the maintenance system includes an OOS-based recovery process, represented by the red arrows in Fig.
+1.
+If an operational satellite experiences a failure mode that is recoverable through OOS, the failed satellite remains near the target orbit of its original slot and waits for service.
+A servicer spacecraft deployed in advance by the OOS provider then performs OOS, after which the failed satellite is retained as an in-plane spare.
 2.2 Supply Chain Overview and Assumptions
-We assume that the target system for maintenance has a Walker Delta pattern [21], with inclination $i$ and altitude $h_{\mathrm{plane}}$.
-It consists of $N_{\mathrm{plane}}$ orbital planes whose right ascensions of ascending nodes (RAANs) are equally spaced by $2\pi/N_{\mathrm{plane}}$.
+The satellite constellation considered in the proposed strategy is assumed to follow a Walker Delta pattern with inclination $i$ and altitude $h_{\mathrm{plane}}$.
+The constellation consists of $N_{\mathrm{plane}}$ orbital planes, whose RAANs are equally spaced by $2\pi/N_{\mathrm{plane}}$.
 Each orbital plane contains $N_{\mathrm{sat}}$ operational satellites.
-The operator sources spare satellites in two ways: 1) recovering the operational life of failed satellites via OOS, and 2) launching newly manufactured satellites from the ground.
-Figure 1 presents an overview of the proposed supply chain.
+As noted earlier, the constellation operator sources spare satellites through two channels: (1) recovering the operational life of failed satellites via OOS and (2) launching newly manufactured satellites from the ground.
+Figure 2 provides an overview of the proposed supply chain.
 Figure 2: Overview of the supply chain for the OOS-integrated maintenance strategy.
@@
 As a proxy for $T_{\mathrm{life}}$, we introduce the maximum number of OOS per satellite, $N_{\mathrm{oos}}$, as a tracking parameter.
-The service lead time, defined as the duration between service request reception and service completion, is primarily composed of a series of maneuvers performed by the servicing spacecraft to approach the target satellite (rendezvous, proximity operations, docking), along with the time required to execute the servicing activity itself.
-The lead time depends on the current state of the OOS provision system.
-For example, if an available servicing spacecraft is spatially distant from the target satellite at the time of the service request, the lead time will be long.
-Since the time required for the available servicing spacecraft to maneuver to the target is inherently random, the lead time is modeled as a probabilistic variable.
+In this study, we consider the aspects most relevant to the customer, namely, service price and responsiveness.
+Specifically, we assume that the OOS provider has already deployed the infrastructure required to support the constellation under consideration, including orbital depots for fuel, replacement units, and other servicing resources, as well as multiple servicing spacecraft capable of approaching and restoring failed satellites.
+Under this assumption, the OOS system has sufficient capacity to respond to service requests with stable cost and performance.
+The OOS price is the amount the constellation operator pays to use the OOS.
+In this study, OOS is charged per servicing event at a fixed price of $p_{\mathrm{oos}}$.
+Correspondingly, the OOS provider receives revenue of $p_{\mathrm{oos}}$ per servicing event.
+To provide OOS, the provider must allocate resources to each servicing event, and we model the corresponding per-event cost as $c_{\mathrm{oos}}$.
+As a result, the OOS provider’s profit per servicing event is $\left(p_{\mathrm{oos}}-c_{\mathrm{oos}}\right)$.
+Service lead time is the representative measure of service responsiveness.
+It is defined as the duration between receipt of a service request and completion of the requested service, and it is the key performance parameter considered in this study.
+It consists primarily of the time required for the sequence of maneuvers by which a servicing spacecraft approaches the target satellite, including rendezvous, proximity operations, and docking, together with the time required to perform the servicing activity itself.
+The lead time depends on the current state of the OOS system.
+For example, if an available servicing spacecraft is far from the target satellite when the request is made, the resulting lead time is longer.
+Because the time required for a servicing spacecraft to reach the target is inherently uncertain, we model service lead time as a random variable.
+The modeling details are presented later in this section.
+Note that the cost ($c_{\mathrm{oos}}$) and the responsiveness of OOS are closely related.
+Given the OOS technologies, service types, and target failure modes of the serviced spacecraft, the OOS provider’s operational effort determines both service responsiveness and cost.
+For example, if the OOS provider allocates more resources, such as additional servicing spacecraft, depots, and propellant for faster orbital transfers, the servicing cost increases, but so does responsiveness.
+This relationship is modeled in a simplified manner in this study and is introduced later in Section 4.
 2.2.2 Sourcing Method 2: Deploying New Satellites
@@
 There are $N_{\mathrm{parking}}$ parking planes, with RAANs equally spaced by $2\pi/N_{\mathrm{parking}}$.
-In-plane spares directly replace failed satellites and are distributed across the (near-)target orbits (i.e., the operational planes of the constellation).
+In-plane spares directly replace failed satellites.
 When the inventory of in-plane spares reaches a reorder point $s$, a batch of $Q$ spares is sourced from the parking spares, the next echelon.
@@
     In-orbit failures of operational satellites follow a Poisson process with constant rate $\lambda_{\mathrm{sat}}$.
+This assumption is reasonable under the steady-state operation of mega-constellations, where individual satellite health variations are averaged out at the fleet level, yielding an approximately constant aggregate failure rate.
   2. 2.
-    Spare satellites’ failures are neglected (i.e., ground, parking, and in‐plane spares are assumed failure-free).
+    Failures of spare satellites are neglected; that is, parking and in-plane spares are assumed to be failure-free.
+This simplification is justified because the time spent in spare status is much shorter than the operational lifetime and can therefore be neglected at the strategic level.
   3. 3.
-    In-plane spares are available instantaneously to replace a failed satellite [4].
+    In-plane spares are assumed to be available instantaneously to replace a failed satellite.
+This is reasonable because the associated response time is short, typically on the order of a few days [16], relative to other relevant time scales such as RAAN alignment, launch lead time, orbital transfer, and OOS lead time.
   4. 4.
-    Each target orbit is replenished from the closest available parking orbit (i.e., a nonempty parking orbit with the earliest RAAN alignment).
-When an in-plane stock level hits its reorder point ($s$), a batch of size $Q$ is drawn from that parking orbit [4].
+    Each target orbit is replenished from the closest available parking orbit, defined as the nonempty parking orbit with the earliest RAAN alignment.
+When the inventory of in-plane spare reaches the reorder point, $s$, a batch of size $Q$ is dispatched from that parking orbit.
   5. 5.
-    Ground-to-parking delivery can supply only a single parking orbit [4].
+    Ground-to-parking delivery is assumed to supply only a single parking orbit.
+Without an orbital transfer vehicle or additional mechanisms, supplying multiple parking orbits would be inefficient [71].
   6. 6.
-    A new in-plane order is placed only after the previous one has completed transit to facilitate the tracking of the orders [4].
+    A new in-plane replenishment order is placed only after the previous order has completed transit to facilitate the tracking of the orders.
   7. 7.
-    Stock-outs are rare; therefore, the probability of all parking orbits being out of stock is negligible [4].
+    Stock-outs are assumed to be rare.
+Accordingly, the probability that all parking orbits are simultaneously out of stock is treated as negligible, consistent with nominal maintenance system operations.
   8. 8.
-    Both the reorder point and order quantity at parking orbits are constrained to integer multiples of the batch size ($Q$) [4].
+    The reorder point and order quantity at parking orbits are restricted to integer multiples of the batch size ($Q$).
   9. 9.
-    An OOS lead time follows an exponential distribution, which is a stochastic distribution characterized by a single parameter, with mean time to recovery (or equivalently, mean time to repair, MTTR), $\mu_{\mathrm{oos}}^{-1}$.
+    The OOS system is assumed to have sufficient capacity to serve the target constellation, so that individual service requests may be treated independently and significant waiting time is rare.
+Under this assumption, OOS lead time is modeled as an exponential random variable.
+This choice is motivated by its common use for lead times with these characteristics and by its mathematical tractability [33, 72].
+The corresponding mean time to recovery, or equivalently mean time to repair (MTTR), is given by $\mu_{\mathrm{oos}}^{-1}$.
+Assumptions 1–8 are adapted from prior works [31, 32].
+These assumptions are appropriate for strategic-level analysis of mega-constellations because they preserve the key operational characteristics while reducing the model’s complexity.
+Assumption 9 is newly introduced in this study.
+In the model validation (Section 3), we relax this assumption to examine whether the proposed model remains robust under alternative lead time conditions.
 2.3 In-Plane Spares
@@
 Figure 3 provides a closer view of the closed-loop system shown in Fig.
-1, illustrating the flows of spare and failed satellites for the case of $N_{\mathrm{oos}}=2$.
-Since our model represents the mean behavior over an infinite time horizon, we define $\gamma_{0}$ as the fraction of newly arrived in-plane spares sourced from parking orbits, and $\gamma_{m}$ ($m=1,2,\dots,N_{\mathrm{oos}}$) as the fraction of in-plane spares that have been serviced $m$ times, also in the infinite time horizon.
+2, illustrating the flows of spare and failed satellites for the case of $N_{\mathrm{oos}}=2$, meaning that a satellite can be recovered at most twice.
+From the perspective of replacing the slot vacated by a failed satellite, a spare newly transferred from the parking orbit and a spare recovered through OOS are operationally equivalent.
+However, for modeling purposes, we distinguish in-plane spares according to their service history, namely whether they are newly arrived from the parking orbit or have been recovered through OOS, and, in the latter case, by the number of times they have been serviced.
+After a parking spare arrives in the target orbit, it remains as an in-plane spare until an operational satellite fails and vacates a slot.
+The spare then fills that slot and operates until it fails.
+If the failure is serviceable, which occurs for a fraction $r_{\mathrm{oos}}$ of failures, the failed satellite waits for OOS and returns as an in-plane spare after the service lead time.
+Otherwise, it is disposed of.
+This process repeats until the satellite experiences a non-serviceable failure or reaches the maximum allowable number of services, $N_{\mathrm{oos}}$.
+Over an infinite horizon, let $\gamma_{0},\gamma_{1},\dots,\gamma_{N_{\mathrm{oos}}}$ denote the steady-state fractions of in-plane spares.
+Here, $\gamma_{0}$ denotes newly arrived spares from parking orbits, while $\gamma_{m}$ ($m=1,2,\dots,N_{\mathrm{oos}}$) denotes spares that have undergone service exactly $m$ times.
+Because in-plane spares wait to replace failed operational satellites, they are held as spare inventory.
+We denote their mean stock level by $\overline{\mathrm{SL}}_{\mathrm{plane}}$, which includes both newly arrived spares from parking orbits and spares that have been serviced any number of times.
+Likewise, serviceable failed satellites are held while awaiting service.
+We denote the corresponding mean stock level for satellites awaiting the $m$th service by $\overline{\mathrm{SL}}_{\mathrm{wait},m}$.
 Figure 3: Example spare/failed satellites flows within a target orbit, $N_{\mathrm{oos}}=2$.
 Assuming that the service lead time follows an exponential distribution with the mean value of $\mu_{\mathrm{oos}}^{-1}$ implies that the OOS process recovers a failed satellite with probability $\mu_{\mathrm{oos}}$ per unit time [72].
-Based on this assumption, the flow balance equations at the steady-state for failed satellites awaiting OOS and in-plane spares over an infinite horizon are as follows:
+The corresponding steady-state flow balance equations over an infinite horizon for 1) failed satellites awaiting the $m$th OOS and 2) in-plane spares that have been serviced $m$ times are indicated in Fig.
+3 by the star pairs labeled 1-2 and 2-3, respectively.
+Within each color group (yellow and green), adjacent star pairs represent individual flow-balance relations, corresponding to Eqs.
+(6) and (7):
 $\lambda_{\mathrm{plane}}\gamma_{m-1}r_{\mathrm{oos}}=\overline{\mathrm{SL}}_{\mathrm{wait},m}\mu_{\mathrm{oos}},\text{ for }m=1,2,\dots,N_{\mathrm{oos}},$ (6)
@@
 $f_{\mathrm{plane}}\left(\tau\right)=P^{\prime}_{\mathrm{av}}\left(j\right)\frac{N_{\mathrm{parking}}\left|\dot{\Omega}_{\mathrm{rel}}\right|}{2\pi},\text{ for }\tau\in\mathcal{T}_{j},\ j=1,2,\dots,N_{\mathrm{parking}},$ (13)
-where $P_{\mathrm{av}}^{\prime}\left(j\right)$ is the normalized probability that the $j$th closest parking orbit is the closest available parking one, $\dot{\Omega}_{\mathrm{rel}}$ is the relative RAAN drift rate between target and parking orbits, and $\mathcal{T}_{j}$ is the admissible range of alignment time between a target orbit and its $j$th closest parking orbit, defined as:
+where $P_{\mathrm{av}}^{\prime}\left(j\right)$ is the normalized probability that the $j$th closest parking orbit is the closest available parking one, $\dot{\Omega}_{\mathrm{rel}}$ is the relative RAAN drift rate between target and parking orbits, and $\mathcal{T}_{j}$ denotes the admissible range of lead times for sourcing from the $j$th closest parking orbit, including both the alignment time between the target orbit and its $j$th closest parking orbit and the orbit-raising time from the parking orbit to the target orbit, defined as:
 $P_{\mathrm{av}}^{\prime}\left(j\right)=\frac{P_{\mathrm{av}}\left(j\right)}{\sum_{j^{\prime}=1}^{N_{\mathrm{parking}}}P_{\mathrm{av}}\left(j^{\prime}\right)},\quad P_{\mathrm{av}}\left(j\right)=\left(1-\rho_{\mathrm{parking}}\right)^{j-1}\rho_{\mathrm{parking}},\text{ for }j=1,2,\dots,N_{\mathrm{parking}},$ (14)
@@
 $\mathcal{T}_{j}=\left[\left(j-1\right)\frac{2\pi}{h_{\mathrm{parking}}}\frac{1}{\left|\dot{\Omega}_{\mathrm{rel}}\right|}+t_{\mathrm{trans}},j\frac{2\pi}{N_{\mathrm{parking}}}\frac{1}{\left|\dot{\Omega}_{\mathrm{rel}}\right|}+t_{\mathrm{trans}}\right),\text{ for }j=1,2,\dots,N_{\mathrm{parking}}.$ (16)
-Here, $t_{\mathrm{trans}}$ is the time of flight from a parking orbit to a target orbit calculated by Eq. (3).
+Here, $t_{\mathrm{trans}}$ is the time of flight from a parking orbit to a target orbit calculated by Eq.
+(3), and the right-open interval of $\mathcal{T}_{j}$ is used to avoid overlap between adjacent admissible intervals, thereby preventing double counting at the boundary.
 Following these demand and supply models, we define several inventory parameters.
@@
 $\mathrm{ES}_{\mathrm{parking}}=\int_{\tau\geq t_{\mathrm{lau}}}\sum_{\delta\geq k_{s}}\left(\delta-k_{s}\right)P_{\mathrm{parking}}\left(\delta;\tau\right)f_{\mathrm{parking}}\left(\tau\right)\,d\tau,$ (27)
-$\rho_{\mathrm{parking}}=1-\frac{\mathrm{ES}_{\mathrm{parking}}}{k_{Q}}.$ (28)
+$\rho_{\mathrm{parking}}=1-\frac{\mathrm{ES}_{\mathrm{parking}}}{k_{Q}}{\color[rgb]{0,0,0}{,}}$ (28)
+where $N_{t}$ is the number of discretized time units per year.
 2.5 Annual Maintenance Cost and Mean Time to Disposal
@@
 $T_{\mathrm{opr}}=\frac{\left(N_{\mathrm{oos}}+1\right)N_{t}}{\lambda_{\mathrm{sat}}},$ (39)
-$T_{\mathrm{oos}}=\frac{N_{\mathrm{oos}}}{\mu_{\mathrm{oos}}}.$ (40)
-2.5 Model Validation
-To evaluate the accuracy of our inventory management model, we compared its parameter estimates to the mean of 100 Monte Carlo simulation results for 25 randomly selected instances.
+$T_{\mathrm{oos}}=\frac{N_{\mathrm{oos}}}{\mu_{\mathrm{oos}}}$ (40)
+Note that $T_{\mathrm{parking}}$, $T_{\mathrm{plane}}$, and $T_{\mathrm{oos}}$ are obtained as the ratios of the corresponding stock levels to their outflow rates, $T_{\mathrm{trans}}$ follows from the definition of the mean of a continuous random variable with probability density function $f_{\mathrm{plane}}$, and $T_{\mathrm{opr}}$ is given by the reciprocal of the satellite failure rate.
+The factors ($N_{\mathrm{oos}}+1$) and $N_{\mathrm{oos}}$ arise because the satellite passes through the in-plane spare and operational phases ($N_{\mathrm{oos}}+1$) times, whereas it undergoes on-orbit servicing only $N_{\mathrm{oos}}$ times before disposal.
+3 Model Validation
+The analytical model for the performance parameters constructed in Section 2 is validated by comparing its values with those obtained from discrete simulations.
+We first describe the generation of the validation instances and then present the results.
+3.1 Instance Generation
+To evaluate the accuracy of the proposed inventory management model, we compared its estimates with the sample means obtained from $100$ Monte Carlo simulation runs.
 Each simulation spans $60$ years and is discretized into one-day intervals.
-The fixed parameters and trade space for randomly sampled parameters are in Tables 2 and 3.
-We repeatedly sampled the varying parameters until we obtained 25 instances satisfying the following conditions:
-$\rho_{\mathrm{plane}}\geq 0.98,$ (41)
-$\rho_{\mathrm{parking}}\geq 0.98,$ (42)
-$T_{d}/N_{t}\leq T_{\mathrm{life}},$ (43)
-$s\leq Q,$ (44)
-$k_{s}\leq k_{Q}.$ (45)
-Here, Eqs. (41) and (42) represent the rare shortage assumptions in Section 2.1; Eq. (43) imposes the lifespan condition; and Eqs. (44) and (45) maintain the inventory’s cyclic behavior.
+Some parameters were held fixed, while others were randomly sampled.
+Table 2 lists the fixed parameters, and Table 3 gives the ranges for the sampled parameters.
+Most parameter values were adopted from previous work [32], whereas the OOS-related parameters were defined in this study to span a broad range of plausible values.
+The OOS price, $p_{\mathrm{oos}}$, was set to \$22M, which is approximately equal to the sum of the satellite production cost, $c_{\mathrm{sat}}=$\$0.50.5M, and the per-satellite launch cost assuming a launch capacity of $40$ satellites, i.e., $c_{\mathrm{lau}}/40\approx$\$1.71.7M.
+The same launch capacity was adopted as the reference launcher specification in the case study, following previous work [32].
+The serviceable failure fraction, $r_{\mathrm{oos}}$, was varied from $0.1$ to $0.9$.
+The OOS MTTR, $\mu_{\mathrm{oos}}^{-1}$, was varied from $1$ week to $3$ months to represent a wide range of plausible service-performance levels in LEO [77].
+The maximum number of services, $N_{\mathrm{oos}}$, was varied from $1$ to $4$, based on the feasible range under the fixed satellite lifetime of $T_{\mathrm{life}}=30$ years.
 Table 2: Fixed parameters for validation study
 Parameter, Unit Notation Value
-Fuel mass conversion coefficient, M\$/kg $\epsilon_{\mathrm{fuel}}$ 0.01
-Satellite production cost, M\$/satellite $c_{\mathrm{sat}}$ 0.5
-Spare/failed satellite holding cost, M\$/satellite/year $c_{\mathrm{hold}}$ 0.5
-Launch cost, M\$/launch $c_{\mathrm{lau}}$ 67
+Fuel mass conversion coefficient, \$M/kg $\epsilon_{\mathrm{fuel}}$ 0.01
+Satellite production cost, \$M/satellite $c_{\mathrm{sat}}$ 0.5
+Spare/failed satellite holding cost, \$M/satellite/year $c_{\mathrm{hold}}$ 0.5
+Launch cost, \$M/launch $c_{\mathrm{lau}}$ 67
 Specific impulse of satellite’s propulsion system, sec $I_{\mathrm{sp}}$ 1200
@@
 Satellite’s dry mass, kg $M_{\mathrm{dry}}$ 150
-OOS price, M\$ $p_{\mathrm{oos}}$ 2
+OOS price, \$M $p_{\mathrm{oos}}$ 2
 Lifespan of satellite, years $T_{\mathrm{life}}$ 30
@@
 Satellite failure rate, failures/year $\lambda_{\mathrm{sat}}$ $\left[0.01,0.2\right]$
-Number of orbital planes in the constellation, planes $h_{\mathrm{plane}}$ $\left[20,40\right]$
+Number of orbital planes in the constellation, planes $N_{\mathrm{plane}}$ $\left[20,40\right]$
 Number of parking orbits, planes $N_{\mathrm{parking}}$ $\left[1,20\right]$
@@
 Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
-The absolute error evaluates rational parameters (e.g., $\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$):
+We repeatedly sampled the varying parameters until we obtained $100$ instances satisfying the following conditions:
+$\rho_{\mathrm{plane}}\geq\rho_{\mathrm{plane}}^{\mathrm{req}},$ (41)
+$\rho_{\mathrm{parking}}\geq\rho_{\mathrm{parking}}^{\mathrm{req}},$ (42)
+$T_{d}/N_{t}\leq T_{\mathrm{life}},$ (43)
+$s\leq Q,$ (44)
+$k_{s}\leq k_{Q}.$ (45)
+Here, Eqs.
+(41) and (42) represent the rare shortage assumption (Assumption 7) in Section 2.2 and are also used for service level requirements in later sections.
+Equation (43) imposes the lifespan condition, and Eqs.
+(44) and (45) maintain the cyclic behavior of the inventory.
+For $\rho_{\mathrm{plane}}^{\mathrm{req}}$ and $\rho_{\mathrm{parking}}^{\mathrm{req}}$, we considered values of $0.98$, $0.95$, and $0.90$, which yielded three sets of $100$ instances for testing stress cases under the rare shortage assumption.
+3.2 Results
+We tracked the mean stock levels of in-plane spares ($\overline{\mathrm{SL}}_{\mathrm{plane}}$), parking spares ($\overline{\mathrm{SL}}_{\mathrm{parking}}$), and failed satellites awaiting service ($\overline{\mathrm{SL}}_{\mathrm{wait}}$); the order frequencies at the in-plane and parking spare echelons ($n_{\mathrm{plane}}$ and $n_{\mathrm{parking}}$); the OOS frequency ($n_{\mathrm{oos}}$); the mean time to disposal for satellites serviced up to the maximum allowable number of times ($T_{d}$); the annual maintenance cost ($\mathrm{AMC}$); and the order fill rates of the in-plane and parking spare echelons ($\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$).
+For the rate parameters, $\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$, we used the absolute error:
 $\left|\mathrm{value}_{\mathrm{sim}}-\mathrm{value}_{\mathrm{model}}\right|,$ (46)
@@
 $\frac{\left|\mathrm{value}_{\mathrm{sim}}-\mathrm{value}_{\mathrm{model}}\right|}{\mathrm{value}_{\mathrm{sim}}}\times 100.$ (47)
-The results in Table 4 indicate that relative errors remain below 2% and absolute errors below 1%p, meaning that the constructed model accurately estimates the parameters.
-Table 4: Averaged errors of result parameters
-Result Notation Mean Error
+To construct stress cases for Assumption 9, which assumes an exponential distribution for the OOS lead time, we performed simulations using alternative distributions for lead time with the same MTTR, $\mu_{\mathrm{oos}}^{-1}$, but different coefficients of variation (CVs).
+Here, CV is defined as the ratio of the standard deviation to the mean.
+The tested cases were deterministic, gamma, and lognormal.
+For the gamma distribution, we considered CV values of $1/2$, $\sqrt{2}, 11$, $\sqrt{2}$, and $2$, whereas for the lognormal distribution, we considered CV values of $\sqrt{2}$, $2$, and $4$.
+The deterministic case corresponds to a fixed lead time equal to $\mu_{\mathrm{oos}}^{-1}$.
+The gamma distribution was included because it generalizes the exponential distribution, with CV =1=1 corresponding to the baseline exponential case (Assumption 9).
+Therefore, both lower- and higher-variability cases relative to this baseline were tested.
+The lognormal distribution was included to examine long-tail stress cases, with CV values exceeding the baseline value of $1$.
+The results are summarized in Tables 46, corresponding to the cases where $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.98$, $0.95$, and $0.90$, respectively.
+Table 4: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.98$
+Deterministic Gamma Lognormal
+$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
 Relative Errors (%)
-Mean stock level at in-plane spare $\overline{\mathrm{SL}}_{\mathrm{plane}}$ 1.95
-Mean stock level at parking spare $\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.74
-Mean stock level of failed satellites waiting for service $\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.85
-Order frequency at in-plane spare $n_{\mathrm{plane}}$ 0.49
-Order frequency at parking spare $n_{\mathrm{parking}}$ 1.88
-Service frequency $n_{\mathrm{oos}}$ 0.66
-Mean time to disposal of satellites serviced up to the maximum $T_{d}$ 0.35
-Annual maintenance cost $\mathrm{AMC}$ 0.71
+$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 1.89 1.93 1.92 1.90 1.96 1.91 1.89 1.86 1.87
+$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.73 0.76 0.75 0.66 0.76 0.70 0.75 0.76 0.82
+$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.08 1.95 2.01 2.06 2.51 2.75 2.01 2.29 3.04
+$n_{\mathrm{plane}}$ 0.54 0.50 0.52 0.56 0.45 0.47 0.55 0.54 0.50
+$n_{\mathrm{parking}}$ 1.51 1.71 1.64 1.40 1.46 1.38 1.54 1.72 1.53
+$n_{\mathrm{oos}}$ 0.64 0.76 0.75 0.70 0.76 0.70 0.66 0.77 0.67
+$T_{d}$ 0.43 0.42 0.46 0.39 0.44 0.46 0.48 0.48 0.48
+$\mathrm{AMC}$ 0.62 0.60 0.60 0.60 0.66 0.59 0.56 0.66 0.70
 Absolute Errors (%p)
-Order fill rate at in-plane spare $\rho_{\mathrm{plane}}$ 0.04
-Order fill rate at parking spare $\rho_{\mathrm{parking}}$ 0.05
+$\rho_{\mathrm{plane}}$ 0.10 0.10 0.10 0.09 0.10 0.11 0.10 0.11 0.10
+$\rho_{\mathrm{parking}}$ 0.07 0.06 0.05 0.05 0.06 0.07 0.07 0.06 0.06
+Table 5: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.95$
+Deterministic Gamma Lognormal
+$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
+Relative Errors (%)
+$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 2.65 2.71 2.71 2.65 2.72 2.72 2.69 2.64 2.67
+$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.84 0.82 0.85 0.81 0.88 0.76 0.87 0.80 0.89
+$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.18 1.83 1.92 2.03 2.35 2.53 1.98 2.21 2.95
+$n_{\mathrm{plane}}$ 0.55 0.51 0.49 0.54 0.44 0.47 0.56 0.51 0.47
+$n_{\mathrm{parking}}$ 1.38 1.51 1.52 1.33 1.39 1.37 1.38 1.63 1.38
+$n_{\mathrm{oos}}$ 0.69 0.78 0.75 0.70 0.77 0.75 0.66 0.77 0.66
+$T_{d}$ 0.44 0.44 0.46 0.42 0.46 0.44 0.47 0.42 0.45
+$\mathrm{AMC}$ 0.71 0.64 0.69 0.63 0.72 0.63 0.64 0.69 0.71
+Absolute Errors (%p)
+$\rho_{\mathrm{plane}}$ 0.34 0.35 0.36 0.33 0.35 0.35 0.35 0.34 0.32
+$\rho_{\mathrm{parking}}$ 0.10 0.10 0.09 0.10 0.09 0.08 0.10 0.09 0.11
+Table 6: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.90$
+Deterministic Gamma Lognormal
+$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
+Relative Errors (%)
+$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 2.74 2.83 2.83 2.73 2.78 2.81 2.80 2.72 2.74
+$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.92 0.83 0.93 0.85 0.97 0.86 0.87 0.84 0.93
+$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.17 1.95 2.13 2.21 2.57 2.74 2.08 2.30 3.19
+$n_{\mathrm{plane}}$ 0.55 0.51 0.56 0.53 0.49 0.52 0.59 0.56 0.53
+$n_{\mathrm{parking}}$ 1.42 1.44 1.57 1.32 1.40 1.27 1.25 1.61 1.29
+$n_{\mathrm{oos}}$ 0.73 0.82 0.79 0.70 0.82 0.76 0.71 0.80 0.72
+$T_{d}$ 0.45 0.46 0.48 0.43 0.49 0.49 0.48 0.45 0.48
+$\mathrm{AMC}$ 0.74 0.67 0.70 0.65 0.71 0.66 0.67 0.71 0.72
+Absolute Errors (%p)
+$\rho_{\mathrm{plane}}$ 0.43 0.44 0.44 0.40 0.42 0.42 0.45 0.41 0.39
+$\rho_{\mathrm{parking}}$ 0.16 0.13 0.14 0.12 0.12 0.13 0.15 0.13 0.13
+The results show that the proposed model remains a good approximation across all tested cases.
+In the baseline case corresponding to Assumption 9, i.e., the gamma distribution with CV =1=1 (equivalently, the exponential distribution), all reported relative errors remain below 33% for the three service level requirements, and both absolute order fill rate errors remain below 0.50.5%p.
+This indicates that the model reproduces the simulation results well under the nominal assumption.
+Among the reported parameters, $\overline{\mathrm{SL}}_{\mathrm{plane}}$ and $\overline{\mathrm{SL}}_{\mathrm{wait}}$ exhibit the largest errors.
+This is expected because these quantities are most directly coupled to the OOS-enabled closed-loop recovery process.
+However, their sensitivities to lead-time variability differ.
+In particular, $\overline{\mathrm{SL}}_{\mathrm{wait}}$ is the most sensitive to the shape of the lead-time distribution: for both the gamma and lognormal families, its error increases as the CV increases, even though the MTTR is held fixed.
+This behavior is consistent with the fact that $\overline{\mathrm{SL}}_{\mathrm{wait}}$ depends most directly on the residence time in the OOS stage.
+By contrast, the error in $\overline{\mathrm{SL}}_{\mathrm{plane}}$ is relatively insensitive to the CV because the in-plane spare stock is influenced not only by recovery through OOS but also by replenishment from parking spares.
+As the required order fill rate is relaxed from $0.98$ to $0.90$, the overall error level increases modestly, but the deterioration remains limited.
+Even in the most severe stress case—a lognormal lead time distribution with CV =4=4 under $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.90$—the largest average relative error is only 3.193.19%, observed for $\overline{\mathrm{SL}}_{\mathrm{wait}}$, while the largest average absolute error in order fill rate across all tested cases is only 0.450.45%p.
+Therefore, although higher-order features of the lead time distribution do affect the prediction of the waiting stock, the resulting loss of accuracy remains small for the tested service-level requirements, including thresholds as low as $0.90$.
+In summary, these results support the use of Assumption 9 as a tractable approximation for system-level modeling.
+Model accuracy is driven primarily by the mean OOS lead time, whereas the exact distributional form plays a secondary role.
+Accordingly, the exponential assumption provides a reasonable balance between analytical simplicity and predictive accuracy for design-space exploration and managerial trade-off analysis, although some caution is warranted when interpreting $\overline{\mathrm{SL}}_{\mathrm{wait}}$ under highly variable, long-tailed lead-time distributions and low service-level requirements.
 4 Decision-Making Scenarios and Optimization Problem Formulations
-Based on the model, we develop decision-making scenarios involving two stakeholders—the constellation operator and the OOS provider—and discuss potential tension between them.
-This section presents the details of the decision-making contexts under consideration and their mathematical formulations as optimization problems.
-3.1 Constellation Operator
-The goal of the constellation operator is to maintain their system at an appropriate service level while minimizing relevant costs.
-Such a decision-making context can be formulated as an optimization problem, $\mathbf{P}_{1}$.
-($\mathbf{P}_{1}$) Minimizing the Annual Maintenance Cost of the Constellation
-$\min\ \mathrm{AMC}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right),$ (48)
+This study considers two primary stakeholders: the constellation operator and the OOS provider.
+The constellation operator aims to maintain the required service level of the constellation while minimizing maintenance cost.
+The OOS provider, in turn, aims to maximize profits by selecting service price and performance levels—in this study, responsiveness—that remain commercially attractive to the constellation operator.
+The decisions of the two stakeholders are strongly coupled: the operator’s maintenance strategy determines the demand for OOS, whereas the provider’s pricing and responsiveness influence the operator’s preferred maintenance strategy.
+Balancing these tensions and designing appropriate incentives to achieve mutually beneficial outcomes that support the broader space industry are important challenges.
+Indeed, the interdependence between satellite operators and OOS providers is sometimes regarded as a chicken-and-egg problem [78].
+We first represent this interaction through the following bi-objective optimization problem, whose components are detailed in this section:
+($\mathbf{P}$) Bi-Objective Optimization for the Constellation Operator and OOS Provider
+$\min\ \mathrm{AMC}\left(\mathbf{x};\mathbf{q}\right),$ (48)
+$\max\ \mathrm{AP}\left(\mathbf{x};\mathbf{q}\right),$ (49)
 subject to
-$\rho_{\mathrm{plane}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{plane}}^{\mathrm{req}}$ (49)
-$\rho_{\mathrm{parking}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{parking}}^{\mathrm{req}}$ (50)
+Eqs. (41)–(45),
 $Qk_{Q}\leq Q_{\max},$ (50)
-Eqs. (43)-(45),
-$\mathbf{x}_{\mathrm{con}}=\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}.$ (52)
-Here, the vector of decision variables, $\mathbf{x}_{\mathrm{con}}$ in Eq.
-(52), includes parameters related to the replenishment policy and the configuration of parking orbits, while $\mathbf{q}_{1}$ contains the remaining parameters.
-Equation (48) defines the objective of the problem (minimizing the annual system maintenance cost).
-Equations (49) and (50) ensure that the maintenance system maintains an appropriate service level.
-Equation (51) constrains the number of loaded satellites so that it does not exceed the launch capacity ($Q_{\max}$).
-Finally, $\mathcal{H}_{\mathrm{parking}}$ denotes the set of available parking orbit altitudes.
-3.2 OOS Provider
-The main goal of an OOS provider is to maximize profit. The annual profits (AP) of the OOS provider are defined as
+$c_{\mathrm{oos}}\left(\mu_{\mathrm{oos}};\mathbf{q}\right)\leq p_{\mathrm{oos}},$ (51)
+$\mathrm{AMC}\left(\mathbf{x};\mathbf{q}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (52)
+$\mathbf{x}=\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}},p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]\in\mathbb{Z}_{++}^{6}\times\mathcal{H}_{\mathrm{parking}}\times\mathbb{R}_{++}^{2}.$ (53)
+Here, $\mathbf{x}$ denotes the unified decision vector, and $\mathbf{q}$ collects all remaining parameters.
+The decision vector contains both the constellation operator’s decision variables, $\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}}\right]$, and the OOS provider’s decision variables, $\left[p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]$.
+In addition, $\mathcal{H}_{\mathrm{parking}}$ denotes the set of available parking orbit altitudes.
+From the constellation operator’s perspective, Eq.
+(48) minimizes the annual maintenance cost.
+The service-level requirements in Eqs.
+(41) and (42) ensure that the maintenance system achieves the required order fill rates for in-plane and parking spares, respectively.
+Equation (50) limits the number of loaded satellites so that it does not exceed the launch capacity, $Q_{\max}$, while Eqs.
+(43)–(45) impose the validity conditions of the analytical model.
+From the OOS provider’s perspective, Eq. (49) represents the maximization of annual profit, defined as
 $\mathrm{AP}=\left(p_{\mathrm{oos}}-c_{\mathrm{oos}}\right)n_{\mathrm{oos}}N_{\mathrm{plane}},$ (54)
-where $p_{\mathrm{oos}}$ and $c_{\mathrm{oos}}$ are the unit price and cost of an OOS service, and $n_{\mathrm{oos}}N_{\mathrm{plane}}$ corresponds to the total annual number of services performed.
-Given the OOS technologies, service types, and target subsystems of the spacecraft to be serviced, the OOS provider’s operational effort determines the service’s responsiveness and cost.
-For example, if the OOS provider allocates more resources (e.g., additional servicing spacecraft, depots, and propellant for faster orbital transfers), the servicing cost increases, but so does responsiveness.
-While a detailed investigation of operational rule-setting and cost estimation should be a separate research topic, here we introduce a first-cut model representing the relationship between cost and responsiveness as follows:
+where $p_{\mathrm{oos}}$ and $c_{\mathrm{oos}}$ are the unit price and cost of an OOS, respectively.
+The term $n_{\mathrm{oos}}N_{\mathrm{plane}}$ represents the total annual number of services performed, computed as the product of the OOS frequency, $n_{\mathrm{oos}}$, and the number of target orbits, $N_{\mathrm{plane}}$.
+Thus, annual profit equals the profit per service multiplied by the total annual service volume.
+Given the OOS technologies, service types, and target subsystems, higher responsiveness generally requires greater operational effort, such as additional servicers, depots, and propellant for faster orbital transfers.
+To capture this trade-off, we adopt the following first-cut model relating cost and responsiveness:
 $c_{\mathrm{oos}}=c_{\mathrm{oos},\min}+\frac{\alpha_{1}}{\left(\mu_{\mathrm{oos}}^{-1}-\mu_{\mathrm{oos},\mathrm{ideal}}^{-1}\right)^{\alpha_{2}}}.$ (55)
-Here, $c_{\mathrm{oos},\min}$ (>0>0) denotes the lower bound of the cost, $\mu_{\mathrm{oos},\mathrm{ideal}}$ (>0>0) represents the upper bound of responsiveness, and $\alpha_{1}$ (>0>0) and $\alpha_{2}$ (>0>0) are shape parameters.
-This function is strictly decreasing with respect to MTTR, exhibiting knee points and two asymptotes parallel to the cost and MTTR axes.
-Figure 3 illustrates the overall trend and the effect of each parameter.
-Such a relationship, where increased effort yields greater outcomes but with a diminishing rate of increment, reflects the trade-off between service cost and responsiveness.
-This behavior is consistent with empirical observations in terrestrial logistics [30] and prior studies on OOS [32, 31, 33].
+Here, $c_{\mathrm{oos},\min}$ (>0>0) denotes the lower bound of the service cost, $\mu_{\mathrm{oos},\mathrm{ideal}}$ (>0>0) denotes the upper bound of service responsiveness, and $\alpha_{1}$ (>0>0) and $\alpha_{2}$ (>0>0) are shape parameters.
+This function is strictly decreasing with respect to MTTR and exhibits diminishing returns as responsiveness approaches its ideal bound.
+Such a relationship, in which greater operational effort yields improved responsiveness at a diminishing rate, is consistent with empirical observations in terrestrial logistics [79] and prior OOS-related studies [50, 77, 51].
 Figure 4: Visualization of the adopted cost-responsiveness function with the effect of each parameter (1: $c_{\mathrm{oos},\min}$, 2: $\mu_{\mathrm{oos},\mathrm{ideal}}$, 3: $\alpha_{1}$, 4: $\alpha_{2}$).
-With the aid of the cost-responsiveness function, we formulate the optimization problem $\mathbf{P}_{2}$, representing the decision-making of the OOS provider:
-($\mathbf{P}_{2}$) Maximizing Profits of OOS Operations
-$\max\ \mathrm{AP}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right),$ (55)
-subject to
-$c_{oos}\left(\mu_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq p_{\mathrm{oos}},$ (56)
-$\mathrm{AMC}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (57)
-$\mathbf{x}_{\mathrm{oos}}=\left[p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]\in\mathbb{R}_{++}^{2}.$ (58)
-Here, the vector of decision variables $\mathbf{x}_{\mathrm{oos}}$ in Eq.
-(58) consists of two components: the price of the OOS per service, $p_{\mathrm{oos}}$, and the service responsiveness, $\mu_{\mathrm{oos}}$.
-Note that $\mu_{\mathrm{oos}}$ is the inverse of the MTTR.
-The vector $\mathbf{q}_{2}$ contains the remaining parameters, including those associated with the maintenance strategy of the constellation operator, which served as decision variables in $\mathbf{P}_{1}$.
-The objective function in Eq.(55) seeks to maximize the profits.
 Equation (51) ensures that the OOS price is no less than the corresponding service cost.
-Equation (56) ensures that the OOS price is no less than the corresponding service cost.
-Equation (57) constrains the OOS provider to set its service charge and responsiveness within a range that prevents the constellation operator from abandoning the OOS-integrated strategy.
-The term $\mathrm{AMC}{\mathrm{ref}}$ serves as a reference maintenance cost for adopting OOS; it can be interpreted as the maintenance cost without OOS or as a value that includes a premium over the reference cost from an alternative, non-OOS strategy.
-This premium reflects the additional complexity and uncertainty introduced by OOS integration, for which the operator may require greater expected benefits.
-3.3 Tension Between Constellation Operator & OOS Provider
-Previous subsections introduced the optimization problems reflecting the decision-making contexts of two main stakeholders: the constellation operator and the OOS provider.
-Note that these two problems are highly interdependent.
-On one hand, if the OOS provider raises the OOS price to increase profitability, the constellation operator would rely less on OOS for system maintenance and instead deploy new satellites.
-On the other hand, if the constellation operator relies more on OOS, the profitability of the OOS provider would improve without changing its decisions.
-Balancing such tensions and designing appropriate incentives to achieve mutually beneficial outcomes that stimulate the overall space industry are crucial challenges.
-Indeed, the interdependence between satellite operators and OOS providers is sometimes regarded as a chicken-and-egg problem [34].
-The following bi-objective optimization problem expresses the tension between the constellation operator and the OOS provider by combining $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$:
-($\mathbf{P}_{3}$) Bi-Objective Optimization for Multiple Stakeholders
-$\min\ \mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (59)
-$\max\ \mathrm{AP}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (60)
-subject to
-$Qk_{Q}\leq Q_{\max},$ (61)
-$c_{\mathrm{oos}}\left(\mu_{\mathrm{oos}};\mathbf{q}_{3}\right)\leq p_{\mathrm{oos}},$ (62)
-$\mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (63)
-Eqs. (43)-(45),
-$\mathbf{x}=\left[\mathbf{x}_{\mathrm{con}},\mathbf{x}_{\mathrm{oos}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}\times\mathbb{R}_{++}^{2}.$ (64)
-Here, the vector of decision variables, $\mathbf{x}$ in Eq.
-(64), is the concatenation of $\mathbf{x}_{\mathrm{con}}$ and $\mathbf{x}_{\mathrm{oos}}$ from $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, respectively.
-Solving this bi-objective optimization problem yields a family of non-dominated solutions—solutions that are not inferior to any other solution with respect to all objective functions [35]—known as the Pareto front [36, 37].
-To obtain the Pareto front, we employ the non-dominated sorting genetic algorithm II (NSGA-II), a metaheuristic based on the genetic algorithm [38].
+Equation (52) constrains the provider to choose the price and responsiveness so that the resulting OOS-integrated maintenance strategy remains acceptable to the constellation operator.
+The reference value $\mathrm{AMC}_{\mathrm{ref}}$ may be interpreted as the maintenance cost of a benchmark strategy without OOS, or as a premium-adjusted reference value reflecting the additional complexity and uncertainty introduced by OOS integration.
+Solving Problem $\mathbf{P}$ yields a family of non-dominated solutions[80], known as the Pareto front [81, 82].
+In this study, the Pareto front is obtained using the non-dominated sorting genetic algorithm II (NSGA-II), a genetic algorithm-based metaheuristic for multi-objective optimization [83].
 5 Case Study
-This section presents a case study of the proposed bi-objective optimization problem ($\mathbf{P}_{3}$).
+This section presents a case study of the proposed bi-objective optimization problem ($\mathbf{P}$).
 The case study demonstrates the applicability of the proposed framework using a realistically sized constellation and representative satellite specifications.
@@
 5.1 Baseline Scenario
-We established the baseline scenario using the constellation’s system parameters, launch service parameters, and OOS parameters, which are summarized in Tables 5, 6, and 7, respectively.
-The reference annual maintenance cost, $\mathrm{AMC}_{\mathrm{ref}}$, was assessed using the strategy proposed by Jakob et al.
-[4], which has the same supply chain as the one proposed here but without incorporating OOS, and uses the same constellation’s system and launch service parameters.
-The resulting optimal solution is presented in Table 8.
-The optimal maintenance cost is \$925.1M satisfying the service level over requirements.
-In addition to the scenario parameters, the decision variables and their range are summarized in Table 9.
+We established the baseline scenario using the constellation’s system parameters, launch service parameters, and OOS parameters, which are summarized in Tables 7, 8, and 9, respectively.
+The decision variables and their corresponding ranges are listed in Table 10.
+The constellation and launch service parameters in Tables 7 and 8 are adopted from prior studies [31, 32] to reflect realistic, real-world-scale systems.
+By contrast, the OOS-related parameters cannot yet be assigned fully realistic values, because the relevant service capabilities and market conditions remain uncertain.
+Instead, they are specified within plausible ranges based on the existing literature, and their effects on the results are further examined in the parametric study by varying their values.
+In particular, the minimum MTTR is set to $2$ weeks (i.e., $\mu_{\mathrm{oos}}=1/2$), which lies within a reasonable range according to prior studies on OOS responsiveness [50, 77, 51].
+The serviceable failure ratio, $r_{\mathrm{oos}}$, is set to 2525%, considering the maximum fraction of failures that could be induced by a single spacecraft subsystem failure, namely the thruster subsystem, as discussed by Saleh and Castet [84].
+The minimum OOS cost is set equal to the satellite production cost, $c_{\mathrm{sat}}$.
+The shape parameters, $\alpha_{1}$ and $\alpha_{2}$, are both set to 1.
 Table 7: Constellation’s system parameters for baseline scenario
@@
 $\alpha_{2}$ 1
+Table 10: Decision variables and their ranges for baseline scenario
+Decision Variable, Unit Notation Range
+Reorder point of in-plane spares, satellites $s$ $\left[1,20\right]$
+Order quantity of in-plane spares, satellites $Q$ $\left[1,Q_{\max}\right]$
+Reorder point of parking spares, - $k_{s}$ $\left[1,20\right]$
+Order quantity of parking spares, - $k_{Q}$ $\left[1,Q_{\max}\right]$
+Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
+Number of parking orbits, planes $N_{\mathrm{parking}}$ $\left[1,20\right]$
+Altitude of parking orbits, km $h_{\mathrm{parking}}$ $\left[500,1000\right]$
+OOS price, \$M $p_{\mathrm{oos}}$ $\left[c_{\mathrm{oos},\min},c_{\mathrm{oos},\min}+5\right]$
+OOS MTTR, weeks $\mu_{\mathrm{oos}}^{-1}$ $\left(\mu_{\mathrm{oos},{\mathrm{ideal}}}^{-1},12\right]$
+[…]
 Table 11: Optimal solution and result parameters of the baseline scenario without OOS
@@
 $\rho_{\mathrm{parking}}$ % 98.3
-Table 9: Decision variables and their ranges for baseline scenario
-Decision Variable, Unit Notation Range
-$\mathbf{x}_{\mathrm{con}}$
-Reorder point of in-plane spares, satellites $s$ $\left[1,20\right]$
-Order quantity of in-plane spares, satellites $Q$ $\left[1,Q_{\max}\right]$
-Reorder point of parking spares, - $k_{s}$ $\left[1,20\right]$
-Order quantity of parking spares, - $k_{Q}$ $\left[1,Q_{\max}\right]$
-Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
-Number of parking orbits, planes $N_{\mathrm{parking}}$ $\left[1,20\right]$
-Altitude of parking orbits, km $h_{\mathrm{parking}}$ $\left[500,1000\right]$
-$\mathbf{x}_{\mathrm{oos}}$
-OOS price, \$M $p_{\mathrm{oos}}$ $\left[c_{\mathrm{oos},\min},c_{\mathrm{oos},\min}+5\right]$
-OOS MTTR, weeks $\mu_{\mathrm{oos}}^{-1}$ $\left(\mu_{\mathrm{oos},{\mathrm{ideal}}}^{-1},12\right]$
+Using these parameter settings, the bi-objective optimization problem ($\mathbf{P}$) was solved using NSGA-II [83], implemented in Python with the pymoo package [85].
+The algorithm was run with a population size of 400 for 200 generations, while all other algorithm parameters were kept at their default values.
+To assess the reliability of the results, we performed three independent runs with different random seeds.
+Because the ranges of the obtained solutions were sufficiently similar across the three runs in both the decision and objective spaces, we retained the solution set generated using the first random seed.
+The obtained optimal solutions are scattered in the objective space in Fig.
+5.
+The $x$-axis and $y$-axis in Fig.
+5 represent the first objective (i.e., the constellation operator’s annual maintenance cost, $\mathrm{AMC}$) and the second objective (i.e., the OOS provider’s annual profit $\mathrm{AP}$), respectively.
+Since the first objective is minimized and the second is maximized, the utopia point is located at the upper-left corner of the figure, marked by a red star.
 Figure 5: Solutions of the baseline scenario scattered in the objective space (color: OOS price).
+The values and ranges of the decision variables, together with the resulting performance measures of the Pareto-optimal solutions, are summarized in Table 12.
+Figure 5 shows that, under the baseline setting, the Pareto-optimal solutions form an almost linear frontier spanning $\mathrm{AMC}=790.6$$925.0$ \$/year and $\mathrm{AP}=0$$134.5$ \$/year.
+This frontier is bounded by the OOS provider’s non-negative profit condition (Eq.
+(51)) and the constellation operator’s annual maintenance cost threshold, $\mathrm{AMC}_{\mathrm{ref}}$ (Eq.
+(52)).
+In addition, the constraints in Eqs.
+(41)–(43) and Eq.
+(50) are effectively binding along the Pareto frontier.
+Although they are not satisfied at equality for every solution, the remaining slack is small and is attributable to the integrality of the decision variables.
+Relative to the benchmark solution without OOS in Table 11, the constellation operator can reduce annual maintenance cost by as much as \$134.5M/year (14.5%) while maintaining the required service levels.
+Moreover, the Pareto-optimal set is generated by an almost invariant maintenance strategy on the constellation operator side: $s=3$, $Q=4$, $k_{s}=6$, $k_{Q}=10$, $N_{\mathrm{parking}}=7$, $h_{\mathrm{parking}}\in[674.3,700.4]$ km, and $N_{\mathrm{oos}}=4$.
+This indicates that, in the baseline environment, the location of a Pareto solution is governed primarily by the OOS provider’s decision rather than by multiple qualitatively different maintenance strategies for the constellation operator.
+Compared with the optimal strategy without OOS, the OOS-integrated solutions reduce annual launch cost from \$536.0M/year to \$402.4M/year and annual manufacturing cost from \$160.0M/year to \$120.1M/year, both reductions of approximately 25%.
+These savings are partially offset by annual OOS expenditures of \$48.1M/year to \$182.4M/year.
+Since $\gamma_{0}=75.1\%$, approximately 24.9% of failures are handled through OOS rather than by newly launched replacements.
+Along the Pareto frontier, the main variation comes from the OOS price $p_{\mathrm{oos}}$, whereas the optimal OOS MTTR remains near its upper bound, $\mu_{\mathrm{oos}}^{-1}=10.6$$12.0$ weeks.
+This implies that, under the assumed cost-responsiveness relationship, the provider maximizes profit by selecting relatively slow but inexpensive service.
+Accordingly, the Pareto frontier is nearly linear, indicating that movement along it primarily redistributes value between the constellation operator and the OOS provider, rather than substantially altering the physical structure of the maintenance system.
 Table 12: Decision variables and resultant parameters of the Pareto-optimal solutions: baseline scenario
@@
 Decision Variables
-$\mathbf{x}_{\mathrm{con}}$
 $s$, satellites 3
@@
 $N_{\mathrm{oos}}$, - 4
-$\mathbf{x}_{\mathrm{oos}}$
 $p_{\mathrm{oos}}$, \$M [0.6, 2.3]
@@
 $\gamma_{0}$, % 75.1
-The obtained optimal solutions are scattered in the objective space in Fig.
-4.
-The $x$-axis and $y$-axis in Fig.
-4 represent the first objective (i.e., the constellation operator’s annual maintenance cost, $\mathrm{AMC}$) and the second objective (i.e., the OOS provider’s annual profit $\mathrm{AP}$), respectively.
-Since the first objective is minimized and the second is maximized, the utopia point is located at the upper-left corner of the figure, marked by a red star.
-The values and ranges of the decision variables, along with the resulting parameters of the optimal solutions, are summarized in Table 10.
-Interestingly, all solutions share identical or very similar values for the individual components of $\mathbf{x}_{\mathrm{con}}$, decision variables for the constellation operator.
-Notably, the optimal value of $N_{\mathrm{oos}}$ is 4, its upper bound, suggesting that once OOS is adopted, performing as many OOS operations as possible is preferable to alternative maintenance policies in this bi-objective optimization framework with the given parameters.
-Consequently, $T_{d}$ approaches the value of $T_{\mathrm{life}}$, 30 years.
-In contrast, the OOS provider’s decision variables vary significantly, especially $p_{\mathrm{oos}}$, determining the position of each solution in the objective space.
-Specifically, the decisions are significantly influenced by the OOS price: as the price increases, the constellation operator’s expenditure rises, while the OOS provider’s profitability improves.
-On the other hand, the MTTR value is at its worst-case (upper-bound) value of 12, minimizing the OOS provider’s cost associated with service operations.
-Compared with the optimal without-OOS strategy presented in Table 8, the launch cost is reduced from \$536M to \$402.4M (approximately 25% reduction), and the manufacturing cost decreases from \$160M to \$120.1M (also about 25% reduction).
-These reductions result from leveraging OOS, represented by $\gamma_{0}=75\%$, indicating that only 75% of failures are handled by newly launched satellites.
-These values remain consistent across all Pareto-optimal solutions since the constellation operator’s maintenance strategy remains unchanged, and the responsiveness of the OOS service varies only within a narrow range.
-Thus, the annual OOS cost, ranging from \$48.1M to \$182.4M, determines the cost-effectiveness of the OOS-integrated strategy.
-The Pareto-optimal solutions form a line with a slope of 1, as both objectives are linear with respect to $p_{\mathrm{oos}}$ and share the same proportional coefficient, $n_{\mathrm{oos}}N_{\mathrm{oos}}$—the total annual number of OOS service calls.
-Overall, once the OOS technologies, service types, and target spacecraft subsystems are defined, the balance between the constellation operator’s expenditure and the OOS provider’s profitability across the non-dominated solutions depends primarily on the service charge.
-In contrast, the constellation operator’s optimal maintenance strategy remains unchanged under varying prices.
 5.2 Parametric Study
-[…]
+We next examine how the Pareto-optimal solutions change when the OOS-related parameters are perturbed around the baseline scenario.
+Specifically, we vary the serviceable-failure ratio $r_{\mathrm{oos}}$, the lower bound of OOS cost $c_{\mathrm{oos},\min}$, the ideal responsiveness parameter $\mu_{\mathrm{oos},\mathrm{ideal}}$, and the two shape parameters $\alpha_{1}$ and $\alpha_{2}$, as summarized in Table 13.
+All optimization runs were performed using the same algorithm settings as in the baseline scenario.
+To avoid repetitive interpretation, the results are discussed in terms of three questions: how each parameter shifts the Pareto frontier, whether it changes the constellation operator’s preferred maintenance strategy, and what mechanism explains the observed change.
 Table 13: Cost-responsiveness function parameters for parametric study instances
@@
 5.2.1 Serviceable Failures Ratio: $r_{\mathrm{oos}}$
-We first investigate the effect of the serviceable-failure ratio ($r_{\mathrm{oos}}$).
-In Instances 1-1 and 1-2, the $r_{\mathrm{oos}}$ values are set as 0.5 and 0.1, respectively.
-The results are shown in Fig.
-5(a) and Table 12.
+Among the tested parameters, the serviceable-failure ratio $r_{\mathrm{oos}}$ has the strongest structural effect, because it directly determines the share of failures that can be covered by the production-based replenishment to the recovery from OOS.
+As shown in Fig.
+6(a) and Table 14, increasing $r_{\mathrm{oos}}$ from $0.1$ to $0.5$ shifts the Pareto frontier markedly toward the utopia point: the minimum $\mathrm{AMC}$ decreases from \$873.4M/year to \$661.9M/year, while the maximum $\mathrm{AP}$ increases from \$51.7M/year to \$263.1M/year.
+This improvement is accompanied by substantial reductions in annual launch and manufacturing costs, from \$482.4M/year to \$283.7M/year and from \$144.0M/year to \$82.6M/year, respectively.
+At the same time, $\gamma_{0}$ decreases from 90.0% to 51.6%, implying that the fraction of failures handled through OOS increases from 10.0% to 48.4%.
 (a) Instance 1-X: variations in $r_{\mathrm{oos}}$.
@@
 Decision Variables
-$\mathbf{x}_{\mathrm{con}}$
 $s$, satellites 3 4
@@
 $N_{\mathrm{oos}}$, - 4 [3, 4]
-$\mathbf{x}_{\mathrm{oos}}$
 $p_{\mathrm{oos}}$, \$M [0.6, 2.3] [0.6, 2.2]
@@
 $\gamma_{0}$, % 51.6 90.0
-From Fig.
-5(a), we observed that increasing and decreasing $r_{\mathrm{oos}}$ shift the Pareto-optimal solutions toward the upper-left and lower-right directions, respectively—corresponding to movements toward and away from the utopia point.
-Moreover, the Pareto fronts of all instances appear nearly parallel.
-This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance, although differences among instances exist, as shown in Table 12.
-Instance 1-1 has $N_{\mathrm{oos}}$ fixed at its upper bound, whereas Instance 1-2 has $N_{\mathrm{oos}}$ values of 3 and 4, slightly detached from the upper bound.
-However, despite the variation in $N_{\mathrm{oos}}$, the $\gamma_{0}$ value of Instance 1-2 remains at 0.9.
-This is because $r_{\mathrm{oos}}=0.1$ is sufficiently small that changes in $N_{\mathrm{oos}}$ between 3 and 4 have only a minimal effect on $\gamma_{0}$ (Eq.
-(9)).
-Interestingly, although the OOS provider’s allowable decision ranges—those of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$—remain almost the same as in the baseline scenario, the ranges of the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, differ significantly.
-This implies that the constellation operator must reconsider its strategy carefully when $r_{\mathrm{oos}}$ changes substantially to maintain non-dominated solutions, even though the allowable responsiveness and OOS charge remain nearly unchanged.
-By examining the cost components and $\gamma_{0}$, we found that the deployment of new satellites decreases as $r_{\mathrm{oos}}$ decreases.
-Specifically, from Instance 1-1 to Instances 0 and 1-2, the values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$ increase from \$283.7M to \$482.4M, from \$82.6M to \$144.0M, and from 51.6% to 90.0%, respectively, a natural outcome of the reduced contribution of OOS services.
+Unlike the other OOS-related parameters, $r_{\mathrm{oos}}$ also changes the constellation operator’s preferred maintenance configuration.
+When $r_{\mathrm{oos}}=0.5$, the optimal value of $N_{\mathrm{oos}}$ remains fixed at its upper bound.
+In contrast, when $r_{\mathrm{oos}}=0.1$, $N_{\mathrm{oos}}$ takes values of 3 or 4, and the optimal stocking decisions move closer to those of the without-OOS benchmark in Table 11: $s$, $Q$, $k_{Q}$, and $N_{\mathrm{parking}}$ coincide with the benchmark values, while $k_{s}$ and $h_{\mathrm{parking}}$ differ only modestly.
+This result indicates that $r_{\mathrm{oos}}$ is the key parameter determining whether OOS functions merely as an economic supplement or becomes a structurally important component of the maintenance architecture.
 5.2.2 Shift Parameters of Cost-Responsiveness Function: $c_{\mathrm{oos},\min}$ & $\mu_{\mathrm{oos},\mathrm{ideal}}$
-Next, we conducted experiments by varying the shift parameters of the cost-responsiveness function in Eq.
-(54), $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$.
+We next examine the shift parameters of the cost-responsiveness function, $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$.
 The resulting solutions are illustrated in Figs.
-5(b) & 5(c), and the detailed values and ranges of the decision variables and result parameters are provided in Table 13.
+6(b) and 6(c), while the detailed values and ranges of the decision variables and outcome measures are reported in Table 15.
+Their main effect is to shift the Pareto frontier without structural change of the constellation operator’s preferred maintenance strategy.
+Across Instances 2-X and 3-X, the operator-side decision variables remain identical to those of the baseline scenario, and the corresponding dependence on OOS also remains unchanged, as reflected by the constant values of $C_{\mathrm{lau}}=\mathdollar 402.4$M/year, $C_{\mathrm{manufac}}=\mathdollar 120.1$M/year, and $\gamma_{0}=75.1\%$.
 Table 15: Parametric study results: variations in $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$
@@
 Decision Variables
-$\mathbf{x}_{\mathrm{con}}$
 $s$, satellites 3
@@
 $N_{\mathrm{oos}}$, - 4
-$\mathbf{x}_{\mathrm{oos}}$
 $p_{\mathrm{oos}}$, \$M [1.1, 2.3] [0.4, 2.3] [0.6, 2.3] [0.6, 2.3]
@@
 $\gamma_{0}$, % 75.1
-We first observed from Fig.
-5(b) that increasing and decreasing $c_{\mathrm{oos},\min}$ shift the Pareto-optimal solutions toward the lower-right and upper-left directions, respectively.
-Figure 5(c) shows that variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have the opposite effect, but its impact is insignificant.
-This is because the range of optimal $\mu_{\mathrm{oos}}$ lies in the lower-right tail of the cost-responsiveness curve shown in Fig.
-3, where translation along the $x$-axis has only a limited effect on $c_{\mathrm{oos}}$.
-Similar to the results of the previous instances, the Pareto fronts of all cases appear nearly parallel.
-This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance.
-In addition, we can see that the constellation operator’s decision variables remain the same as those in the baseline scenario.
-However, the minimum values of the optimal $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ shift in Instance 2-1 and Instance 2-2, respectively, due to the translation of the cost-responsiveness curve along the $y$- and $x$-axes.
-Because the constellation operator’s decision variables remain identical to those in the baseline scenario, the dependence on OOS does not change, as reflected by the unchanged values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$.
-Nevertheless, variations in service charge and responsiveness lead to changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, through slight adjustments in $C_{\mathrm{hold}}$ and more substantial changes in $C_{\mathrm{oos}}$.
+The effect of $c_{\mathrm{oos},\min}$ is direct and substantial.
+Increasing $c_{\mathrm{oos},\min}$ from \$0.25M to \$1M shifts the Pareto frontier downward and to the right, increasing the minimum $\mathrm{AMC}$ from \$770.7M/year to \$830.5M/year and reducing the maximum $\mathrm{AP}$ from \$154.4M/year to \$94.6M/year.
+By contrast, varying $\mu_{\mathrm{oos},\mathrm{ideal}}$ has only a minor effect: changing it from $0.25$ to $1$ week-1 changes the minimum $\mathrm{AMC}$ only from \$792.7M/year to \$789.9M/year and the maximum $\mathrm{AP}$ only from \$132.4M/year to \$135.2M/year.
+This limited sensitivity arises because the efficient solutions remain in the low-responsiveness region of the cost-responsiveness curve, where horizontal shifts of the asymptote have only a small influence on cost.
 5.2.3 Shape Parameters of Cost-Responsiveness Function: $\alpha_{1}$ & $\alpha_{2}$
-Lastly, we vary the shape parameters of the cost-responsiveness function, $\alpha_{1}$ and $\alpha_{2}$. The resulting solutions are presented in Figs. 5(d) & 5(e) and Table 14.
+Lastly, we vary the shape parameters $\alpha_{1}$ and $\alpha_{2}$, which reshape the curvature of the cost-responsiveness function rather than shift it rigidly.
+The resulting solutions are presented in Figs.
+6(d) & 6(e) and Table 16.
+Their effects are moderate and, again, do not alter the constellation operator’s maintenance strategy.
+Across all four instances, the operator-side decision variables remain identical to those in the baseline scenario, and the values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$ therefore remain unchanged.
 Table 16: Parametric study results: variations in $\alpha_{1}$ and $\alpha_{2}$
@@
 Decision Variables
-$\mathbf{x}_{\mathrm{con}}$
 $s$, satellites 3
@@
 $N_{\mathrm{oos}}$, - 4
-$\mathbf{x}_{\mathrm{oos}}$
 $p_{\mathrm{oos}}$, \$M [0.7, 2.3] [0.6, 2.3] [0.5, 2.3] [0.8, 2.3]
@@
 $\gamma_{0}$, % 75.1
-From Figs.
-5(d) & 5(e), we observed that an increase in $\alpha_{1}$ shifts the Pareto-optimal solutions toward the lower-right direction, whereas a change in $\alpha_{2}$ produces the opposite shift.
-The trend associated with $\alpha_{1}$ is intuitive, as increasing $\alpha_{1}$ raises the OOS cost for a given $\mu_{\mathrm{oos}}$.
-However, the effect of $\alpha_{2}$ depends on the value of $\mu_{\mathrm{oos}}$: when $\mu_{\mathrm{oos}}$ lies near the lower-right tail of the cost-responsiveness curve (i.e., for small $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ increases the OOS cost; conversely, when $\mu_{\mathrm{oos}}$ lies near the upper-left tail (i.e., for large $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ decreases the OOS cost.
-Because the current Pareto-optimal solutions occupy the region near the lower-right tail of the cost-responsiveness curve, variations in $\alpha_{2}$ move the overall solutions in the opposite direction of $\alpha_{1}$.
-As with the previous instances, the Pareto fronts across all cases remain nearly parallel, for the same underlying reasons discussed earlier.
-Notably, the minimum values of $p_{\mathrm{oos}}$ vary between instances, whereas the maximum values remain almost constant, consistent with the previously observed patterns.
-In contrast to earlier cases, the maximum values of $\mu_{\mathrm{oos}}^{-1}$ change when either $\alpha_{1}$ decreases or $\alpha_{2}$ increases, as these adjustments make the cost-responsiveness function more favorable in shape, allowing greater responsiveness to be achieved with only a marginal increase in OOS cost.
-Consistent with Instances 2-X and 3-X, the constellation operator’s decision variables are identical to those in the baseline scenario.
-As a result, the level of dependence on OOS remains unchanged, as reflected in the constant values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$.
-However, variations in the service charge and responsiveness still lead to corresponding changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$.
-4.3 Results Summary and Strategic Implications
-Through the case study, we observed Pareto-efficient solutions of $\mathbf{P}_{3}$ with variations in OOS-related parameters ($r_{\mathrm{oos}}$, $c_{\mathrm{oos},\min}$, $\mu_{\mathrm{oos},\mathrm{ideal}}$, $\alpha_{1}$, $\alpha_{2}$), while keeping the constellation and launch service parameters fixed.
-The summary of the results and their implications is provided below.
-Given the constellation and launch service parameters, as well as the fraction of serviceable failures ($r_{\mathrm{oos}}$), Pareto-optimal solutions share a common maintenance strategy adopted by the constellation operator.
-This observation suggests that, to remain Pareto-efficient, there exists a dominant maintenance strategy for the satellite constellation.
-However, each Pareto-efficient solution differs in the OOS provider’s decision variables, particularly the OOS price $p_{\mathrm{oos}}$, which governs the location of the Pareto-optimal solution in the objective space.
-As $p_{\mathrm{oos}}$ increases, the constellation operator’s annual maintenance cost ($\mathrm{AMC}$) rises, while the OOS provider’s annual profit ($\mathrm{AP}$) also increases.
-Therefore, $p_{\mathrm{oos}}$ acts as a balancing parameter between the constellation operator’s expenditure and the OOS provider’s profits.
-Although the responsiveness of OOS ($\mu_{\mathrm{oos}}$) varies among different Pareto-optimal solutions, the variations remain within a limited range and have only minor effects on the results, since they lie within the right-bottom tail of the cost-responsiveness curve shown in Fig.
-3.
-Consequently, the Pareto-optimal solutions form a linear trend with a slope of 1, cut by lines of $\mathrm{AP}=0$ (as the OOS cost cannot exceed the OOS price) and $\mathrm{AMC}=\mathrm{AMC}_{\mathrm{ref}}$ (due to the maintenance cost threshold set by the constellation operator).
-The slope of 1 is attributed to variations in $p_{\mathrm{oos}}$ affecting $\mathrm{AMC}$ linearly via Eq.
-(34) and $\mathrm{AP}$ linearly via Eq.
-(53), both with coefficients of $n_{\mathrm{oos}}N_{\mathrm{plane}}$.
-Among the OOS-related parameters, the only factor that alters the constellation operator’s maintenance strategy is $r_{\mathrm{oos}}$, the serviceable failure ratio.
-When $r_{\mathrm{oos}}$ is sufficiently high, Pareto-optimal solutions consistently yield the maximum value of $N_{\mathrm{oos}}$, indicating that the constellation operator fully utilizes the OOS capability.
-However, as $r_{\mathrm{oos}}$ decreases significantly, $N_{\mathrm{oos}}$ may deviate from its upper bound, since it has little impact on the fraction of failures replaced by newly launched satellites, represented by the ratio $\gamma_{0}$ (or equivalently, the fraction of failed satellites serviced by OOS, $1-\gamma_{0}$).
-As a result, increasing $r_{\mathrm{oos}}$ raises the fraction of failures handled by OOS, thereby increasing the proportion of the total annual maintenance cost attributed to OOS.
-Conversely, when $r_{\mathrm{oos}}$ decreases, the opposite trend occurs.
-Regarding the cost-responsiveness function, variations in the shift parameters—the ideal OOS cost ($c_{\mathrm{oos},\min}$) and responsiveness ($\mu_{\mathrm{oos},\mathrm{ideal}}$)—minimally affect the decision variables of the Pareto-optimal solutions.
-Since the Pareto-optimal solutions lie within the right-bottom tail of the cost-responsiveness curve, changes in $c_{\mathrm{oos},\min}$ directly shift the OOS cost, whereas variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have negligible influence on the overall cost and results.
-In contrast, variations in the shape parameters $\alpha_{1}$ and $\alpha_{2}$ significantly affect the efficient ranges of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ by reshaping the overall curvature of the cost-responsiveness curve.
-Increasing $\alpha_{1}$ raises the overall cost for a given responsiveness, shifting the Pareto-optimal solutions toward the degrading direction.
-Conversely, because the Pareto-optimal solutions are located near the right-bottom tail of the curve, increasing $\alpha_{2}$ moves the Pareto-optimal solutions toward the improving direction.
+Increasing $\alpha_{1}$ makes the cost-responsiveness curve less favorable and shifts the Pareto frontier away from the utopia point: the minimum $\mathrm{AMC}$ increases from \$786.1M/year for $\alpha_{1}=0.5$ to \$798.6M/year for $\alpha_{1}=2$, while the maximum $\mathrm{AP}$ decreases from \$139.0M/year to \$126.5M/year.
+The effect of $\alpha_{2}$ is reversed in the present problem because the efficient solutions lie near the low-responsiveness tail of the curve.
+Specifically, increasing $\alpha_{2}$ from $0.5$ to $2$ improves the frontier, reducing the minimum $\mathrm{AMC}$ from \$807.9M/year to \$781.8M/year and increasing the maximum $\mathrm{AP}$ from \$117.2M/year to \$143.2M/year.
+Hence, the curvature of the provider’s cost technology matters, but mainly through the economic terms of OOS rather than through a redesign of the constellation maintenance system.
+5.3 Discussions
+5.3.1 Main Findings and Strategic Implications
+Three main findings emerge from the case study.
+First, under fixed constellation, launch service, and OOS-related parameters, the Pareto-optimal set within a given scenario (with fixed OOS-related parameters) is generated by an almost invariant maintenance strategy on the constellation operator’s side.
+In the baseline scenario, the optimal operator-side decisions remain unchanged across the Pareto frontier, while the main variation comes from the provider-side pricing decision, with only limited variation in OOS responsiveness.
+As a result, the baseline Pareto frontier spans approximately $(\mathrm{AMC},\mathrm{AP})=(\mathdollar 925.0\text{M/year},0)$ to $(\mathdollar 790.6\text{M/year},\mathdollar 134.5\text{M/year})$.
+This indicates that, relative to the benchmark solution without OOS, OOS can create up to \$134.5M/year of maintenance-cost savings for the constellation operator while still allowing non-negative profit for the OOS provider.
+A similar within-instance pattern is also observed across every scenario in the parametric study.
+Second, OOS provides material logistical relief even before considering how value is divided between the two stakeholders.
+Relative to the benchmark solution without OOS, the baseline OOS-integrated solutions reduce annual launch cost from \$536.0M/year to \$402.4M/year and annual manufacturing cost from \$160.0M/year to \$120.1M/year, corresponding to reductions of approximately 25% in both components, while maintaining the required service levels of approximately 98% throughout the case study.
+In the baseline case, $\gamma_{0}=75.1\%$, implying that approximately 24.9% of failures are handled through OOS rather than by newly launched replacements.
+This shows that the value of OOS lies not only in the financial transfer between stakeholders but also in its ability to reduce dependence on new satellite production and launch.
+In a broader sense, this may also contribute to more sustainable use of the space environment, although such implications lie beyond the scope of the present model.
+Third, among the OOS-related parameters, the serviceable-failure ratio $r_{\mathrm{oos}}$ is the only one that materially changes the constellation operator’s preferred maintenance architecture.
+Increasing $r_{\mathrm{oos}}$ from 0.1 to 0.5 reduces the minimum $\mathrm{AMC}$ from \$873.4M/year to \$661.9M/year and increases the maximum $\mathrm{AP}$ from \$51.7M/year to \$263.1M/year, while also shifting the operator-side decisions farther from the benchmark solution without OOS and toward stronger reliance on OOS.
+By contrast, variations in $c_{\mathrm{oos},\min}$, $\mu_{\mathrm{oos},\mathrm{ideal}}$, $\alpha_{1}$, and $\alpha_{2}$ mainly shift or reshape the Pareto frontier without materially changing the operator-side configuration.
+From a managerial standpoint, this implies that expanding the set of recoverable failure modes may be at least as important as reducing service cost or improving nominal responsiveness.
+5.3.2 Limitations
+These findings should be interpreted in light of several limitations.
+First, the cost-responsiveness relationship is imposed exogenously through a specific parametric functional form.
+Although its qualitative shape is broadly plausible, given the underlying effort/cost trade-off and its consistency with prior work, it may not accurately represent future OOS technologies, operational concepts, or market structures.
+Second, recovered satellites are assumed to return to an “as-good-as-new” state, which likely overstates the long-run benefit of repeated servicing.
+Third, the analysis adopts a two-stakeholder bi-objective optimization framework and therefore abstracts from bargaining, competition among OOS providers, contracting frictions, and other strategic effects that may determine which solution is realized in practice.
+As a result, the realized outcome need not coincide with the Pareto-optimal frontier identified by the present model and may, in practice, lie outside it.
 6 Conclusions
@@
 To capture the OOS provider’s decision-making context, we further introduce a functional relationship between OOS cost and responsiveness, characterized by a strictly decreasing curve with a knee point and two asymptotes parallel to the coordinate axes, implying cost and performance bounds.
-Finally, we formulate each stakeholder’s problem as an individual optimization model, and then integrate them into a bi-objective optimization framework to identify the family of Pareto-optimal (ideal) solutions.
-Finally, the developed models and optimization framework are applied to realistic large-scale constellation maintenance scenarios under various OOS-related parameters.
-There are several promising avenues for future research.
-First, this study constrains the cost-responsiveness relationship to a specific functional form characterized by two shift and two shape parameters.
-Although this form can be applied to many scenarios, the resulting insights are not universal.
-Moreover, since these four parameters are defined based on current estimates, and OOS technology is still emerging, the relationship between cost and performance remains uncertain and likely to evolve.
-Therefore, further investigation into this relationship would enhance the relevance and generality of the results presented here.
-Second, we assume that a recovered system returns to an “as-good-as-new” state.
-In practice, however, recovered satellites are likely to experience some degree of degradation.
-This factor could affect the effectiveness of OOS as part of the overall maintenance strategy, making it an important topic for future study.
-Third, future research may also explore more detailed and realistic decision-making contexts and relevant solution concepts.
-Although the current solution concept, Pareto optimality, provides valuable theoretical insights, such idealized solutions are not always directly attainable in practice.
-Alternative frameworks may better represent real-world interactions.
-For instance, in situations where the constellation operator and OOS providers act in self-interest, or where multiple OOS providers compete for service opportunities, game-theoretic approaches such as the Nash equilibrium could yield more practical and policy-relevant insights.
+We then formulate the interaction between the two stakeholders—the constellation operator and the OOS provider—as a bi-objective optimization problem to identify the family of Pareto-optimal solutions.
+The resulting models and optimization framework are applied to realistic large-scale constellation maintenance scenarios under various OOS-related parameter settings.
+The case study yields three main findings.
+First, within a given scenario, the Pareto-optimal set is generally generated by an almost invariant maintenance strategy on the constellation operator’s side, while most variation along the Pareto frontier is driven by the provider-side pricing decision.
+Second, integrating OOS improves the maintenance system without degrading the required service levels, indicating that its value lies not only in the economic exchange between stakeholders but also in reducing dependence on conventional replacement through new satellite production and launch.
+Third, among the OOS-related parameters, the fraction of serviceable failure plays the most important structural role, as it is the only parameter that materially changes the constellation operator’s preferred maintenance architecture, while the remaining parameters mainly shift or reshape the Pareto frontier without substantially altering the operator-side decisions.
+Future work should therefore proceed in three main directions.
+First, the cost-responsiveness relationship should be refined using more realistic, technology- and mission-specific models, especially as OOS-related technologies, policies, and market conditions continue to evolve.
+Second, the recovery process could be extended to account for imperfect restoration and state-dependent degradation.
+Third, the present bi-objective framework should be expanded to incorporate more realistic strategic interactions, such as bargaining and competition among multiple OOS providers.
 Acknowledgments